### v1

# On the contraction properties of Sinkhorn semigroups

Abstract
We develop a novel semigroup contraction analysis based on Lyapunov techniques to prove the exponential convergence of Sinkhorn equations on weighted Banach spaces. This operator-theoretic framework yields exponential decays of Sinkhorn iterates towards Schrödinger bridges with respect to general classes of $\phi$-divergences as well as in weighted Banach spaces. To the best of our knowledge, these are the first results of this type in the literature on entropic transport and the Sinkhorn algorithm. We also illustrate the impact of these results in the context of multivariate linear Gaussian models as well as statistical finite mixture models including Gaussian-kernel density estimation of complex data distributions arising in generative models.
Keywords: Entropic Optimal Transport, Sinkhorn semigroups, Schrödinger bridges, contraction coefficients, Lyapunov functions.
Mathematics Subject Classification: 37M25, 49Q22, 47H09, 60J20; secondary: 37M25, 60J05, 94A17.
1 Introduction
1.1 Background
The Schrödinger bridge problem is to find a random evolution with minimal relative entropy w.r.t. some reference physical model and with prescribed distributions of the initial and final time random states. This convex minimization problem is a regularized version of the Monge-Kantorovich optimal transport with a relative entropy constraint, see  [56, 66, 68] for recent surveys with numerous references on this subject.
These optimal transport methodologies have become state-of-the-art tools in a variety of application domains, including generative modeling and machine learning [5, 31, 54, 68], statistical barycenter problems [1, 4, 10, 24, 29, 32], economy [13], computer vision [42, 76], control theory [18, 19], and many others.
Under mild integrability conditions, the Schrödinger bridge problem has a unique solution which can be expressed as the fixed point of Sinkhorn equations also referred to as the iterative proportional fitting procedure. For a more thorough discussion on the origins of Sinkhorn iterations, we refer to the early and pioneering articles [39, 46, 55, 74, 75] as well as the more recent studies [18, 28, 57].
Sinkhorn iterations are based on nonlinear integral transformations over the whole state space. Exact finite-dimensional solutions cannot be computed for general models. However, for finite state spaces the Sinkhorn algorithm is a well known simple procedure that repeatedly applies column and row normalization on matrix spaces [39, 46, 55, 74, 75]. For Gaussian models, Sinkhorn equations can also be solved analytically using sequentially Bayes conjugacy principles [2, 24, 25, 51].
Most of the literature on the exponential convergence of Sinkhorn equations is concerned with finite state spaces [14, 45, 74, 75, 77] as well as compact state spaces or bounded cost functions using Hilbert projective metrics techniques [18, 33, 47, 40]. The recent article [2] also presents a refined convergence analysis for Gaussian models based on matrix Riccati difference equations. It is clearly out of the scope of this article to review all the contributions in this field and we simply refer to the recent book [68] as well as the more recent articles [2, 20, 23] and the references therein.
Sub-linear convergence rates of Sinkhorn equations have also been developed in the articles [3, 17, 44]. Linear rates on non-necessarily compact spaces were first obtained by Léger in [58] using elegant gradient descent and Bregman divergence techniques (see also the recent articles [31, 52]). Refined convergence rates at least one order faster for a general class of models including sub-Gaussian target marginals has also been developed in [48] under some exponential moments estimates on Schrödinger potentials.
There are very few articles on the exponential convergence of Sinkhorn iterates for general models on non-compact spaces and unbounded cost functions that apply to Gaussian models, with the notable exception of two recent significant contributions [23, 20]. In [23], the authors investigate quantitative contraction rates for target marginal distributions with an asymptotically positive log-concavity profile and quadratic cost functions associated with a sufficiently large regularization parameter. These exponential entropic decays were recently refined in [20] to apply to all values of the regularization parameter.
To the best of our knowledge, the exponential convergence of Sinkhorn iterates in the context of statistical mixture models remains an important open research question.
As underlined in [2], Sinkhorn semigroups belong to the class of time varying Markov processes sharing a common invariant target measure at each individual step. These classes of time varying Markov semigroups arise in a variety of areas in applied mathematics including on nonlinear filtering, physics and molecular chemistry, see for instance [6, 34, 35, 36, 70, 71, 72] and references therein.
The stability analysis of time inhomogenous Markov processes is a broad and notoriously difficult subject. The articles [43, 70, 71] provide several successful probabilistic and analytical tools, including coupling methods, spectral analysis as well as functional inequalities such as Nash or log-Sobolev inequalities. Sinkhorn semigroups being defined sequentially by nonlinear conjugate transformations, the application of these refined analytical techniques in this context is rather challenging. There is a need for new tools in order to handle these models.
1.2 Contribution
In this paper, we introduce a novel approach to Sinkhorn exponential convergence analysis that relies on Lyapunov techniques and contraction coefficients on weighted Banach spaces.
To be specific, we investigate Lyapunov approaches and related weighted-norm contraction techniques to analyze the convergence of Sinkhorn semigroups. These semigroup techniques are basic quantitative methods for studying the stability of nonlinear and time-varying stochastic models, see for instance [36] as well as Section 2 in [6]. In the context of time-homogeneous models, we also refer to the pioneering articles by Hasminskii [49] and Meyn and Tweedie [60], see also the book [61] and references therein. These powerful off-the-shelf semigroup techniques are mainly based on two conditions:
  • 1)
    the existence of a Lyapunov function satisfying a drift-type condition, and
  • 2)
    a local minorization condition on the sub-level sets of the Lyapunov function.
Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function. In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator.
Based on these techniques, the analysis developed in this article yields novel contraction estimates with respect to general divergence criteria. It also applies to multivariate Gaussian models as well as statistical finite mixture models.
1.3 Outline of the paper
After presenting some basic notation in Section 1.4, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [2] for a more detailed discussion on these models within our framework). Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C. Section 3.1 recalls some basics facts on the Hilbert projective metric, operator norms, entropy and related $\phi$-divergences. The operator contraction coefficients associated with these criteria are introduced in Section 3.2.
With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups. Section 4.1 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups. The main contraction theorems introduced in this article are presented in Section 4.2 and Section 4.3.
We illustrate the impact of these results in Section 5, which provides several examples, including strong mixing models, Gaussian models as well as Schödinger bridges between finite mixture distributions.
Strong mixing Sinkhorn semigroups are discussed in Section 5.1 in terms of Dobrushin contraction coefficients. In this context, we show that Sinkhorn semigroups contract for any $\phi$-divergence criteria, including the relative entropy, the total variation distance, as well as $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, Jeffreys and Rényi divergences.
Section 5.2 discusses the contraction properties of Gaussian Sinkhorn models. Last but not least, the convergence analysis of Sinkhorn semigroups in the context of statistical finite mixture models, including Gaussian mixture, is carried out in Section 5.3.
Most of the technical proofs, as well as some extended analyses, are provided in Appendices A through E.
1.4 Some basic notation
We let ${\cal B}(\mathbb{X})$ be the algebra of locally bounded measurable functions on a locally compact Polish space $\mathbb{X}$. We denote by ${\cal B}_{b}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ the sub-algebra of bounded measurable functions $f$ on $\mathbb{X}$ endowed with the supremum norm $\|f\|:=\sup_{x\in\mathbb{X}}|f(x)|$. For a given uniformly positive function $\varphi\in{\cal B}(\mathbb{X})$, we let ${\cal B}_{\varphi}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the Banach space of functions $f\in{\cal B}(\mathbb{X})$ equipped with the norm $\|f\|_{\varphi}:=\|f/\varphi\|$.
Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures. Let ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the convex subset of probability measures on $\mathbb{X}$. The Lebesgue integral of a integrable function $f\in{\cal B}(\mathbb{X})$ with respect to (w.r.t.) some $\nu\in{\cal M}(\mathbb{X})$ is denoted by
$\nu(f):=\int_{\mathbb{X}}f(x)~{}\nu(dx)$
For indicator functions $f=1_{A}$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$.
Consider a bounded integral operator ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\cal M}(\mathbb{X})$ and a function $f\in{\cal B}(\mathbb{Y})$. We denote by $\mu{\cal K}\in{\cal M}(\mathbb{Y})$ and ${\cal K}(f)\in{\cal B}(\mathbb{X})$ the measure and the function defined by
$(\mu{\cal K})(dy):=\int~{}\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y)$ (1)
For indicator functions $f=1_{A}$ sometimes we write ${\cal K}(x,A)$ instead of ${\cal K}(1_{A})(x)$. For bounded integral operator ${\cal K}_{1}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and ${\cal K}_{2}(y,dx)$ from $\mathbb{Y}$ into $\mathbb{X}$, we also denote by $({\cal K}_{1}{\cal K}_{2})$ the bounded integral operator from $\mathbb{X}$ into $\mathbb{X}$ defined by the integral composition
$({\cal K}_{1}{\cal K}_{2})(x_{1},dx_{2}):=\int{\cal K}_{1}(x_{1},dy){\cal K}_{2}(y,dx_{2}).$
Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we see that
$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(x,y))=\mu(dy)\,{\cal K}(y,dx).$
A measure $\nu_{1}\in{\cal M}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\cal M}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$. Whenever $\nu_{1}\ll\nu_{2}$, we denote by ${d\nu_{1}}/{d\nu_{2}}$ the Radon-Nikodym derivative of $\nu_{1}$ w.r.t. $\nu_{2}$. We also consider the partial order, denoted $\nu_{1}\leq\nu_{2}$, when $\nu_{1}(A)\leq\nu_{2}(A)$ for any measurable subset $A\subset\mathbb{X}$. The measures are said comparable when we have $c~{}\nu_{1}\leq\nu_{2}\leq c^{-1}~{}\nu_{1}$ for some scalar constant $c>0$.
2 Description of the models
2.1 Entropic optimal transport
Consider locally compact Polish spaces $(\mathbb{X},\lambda)$ and $(\mathbb{Y},\nu)$ endowed with some locally bounded positive measures $\lambda$ and $\nu$. Consider some probability measures of the form
$\lambda_{U}(dx):=e^{-U(x)}~{}\lambda(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy):=e^{-V(y)}~{}\nu(dy)$
for some normalized potential functions $U:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$. Let $\Pi_{U,V}$ be the set of coupling probability measures ${\cal P}$ on the product space $(\mathbb{X}\times\mathbb{Y})$, with prescribed first and second coordinate marginals $(\lambda_{U},\nu_{V})$. Denote ${\cal Q}$ a Markov transition from $\mathbb{X}$ into $\mathbb{Y}$ of the form
${\cal Q}(x,dy)=e^{-W(x,y)}~{}\nu(dy)\quad\mbox{\rm and set}\quad{\cal R}(y,dx):=e^{-W(x,y)}~{}\lambda(dx).$ (2)
In the latter expression, $W(x,y)$ stands for some potential function on $(\mathbb{X}\times\mathbb{Y})$. In optimal transport literature $W$ is sometimes called the cost function.
Consider the reference probability measure
${\cal P}_{0}(d(x,y)):=(\lambda_{U}\times{\cal Q})(d(x,y)):=\lambda_{U}(dx)~{}{\cal Q}(x,dy).$
The entropic transport problem (a.k.a. dynamic Schrödinger bridge problem) w.r.t. the reference measure ${\cal P}_{0}$ is formulated as
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0}),$ (3)
where $\text{Ent}({\cal P}~{}|~{}{\cal P}_{0})=\int\log\frac{\delta{\cal P}}{d{\cal P}_{0}}d{\cal P}$ is the relative entropy (or Kullback-Leibler divergence) between ${\cal P}$ and ${\cal P}_{0}$. It is implicitly assumed there exists some ${\cal P}\in\Pi_{U,V}$ such that $\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{0})<\infty$. This condition ensures the existence of a Schrödinger bridge distribution $\mathbb{P}$ solving (3) (cf. the seminal article by Csiszár [27], as well as Section 6 in the Lecture Notes by Nutz [66], see also the survey article by Léonard [56] and references therein).
2.2 Schrödinger bridges
The Schrödinger bridge distribution $\mathbb{P}$ defined in (3) can rarely be computed analytically. However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [28, 74, 75]. This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions
${\cal P}_{2n}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad{\cal P}_{2n+1}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}$ (4)
starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dx)$ and a transition operator ${\cal S}(x,dy)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
$(\mu\times{\cal S})^{\flat}(d(x,y)):=\mu(dy){\cal S}(y,dx).$
The Markov transitions ${\cal S}_{n}$ in (4) are defined by the Sinkhorn iterations
$\left\{\begin{array}[]{l}(\pi_{2n}\times{\cal S}_{2n+1})^{\flat}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\cal S}_{2(n+1)}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\cal S}_{2n+1}.\end{array}\right.$ (5)
In the dual formulation, the bridge distributions ${\cal P}_{n}$ are often written as
${\cal P}_{n}(d(x,y))=e^{-W(x,y)}~{}\lambda_{U_{n}}(dx)~{}\nu_{V_{n}}(dy),$ (6)
with the initial potential functions $(U_{0},V_{0})=(U,0)$ and potential functions $(U_{n},V_{n})$ defined sequentially by the integral recursions
$U_{2n+1}$ $=$ $~{}U_{2n}~{}:=U+\log{{\cal Q}(e^{-V_{2n}})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2(n+1)}}$ $=$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2n+1}}:=V+\log{{\cal R}(e^{-U_{2n}})}$ (7)
When $n\to\infty$, the sequence of distributions ${\cal P}_{2n}$ converges towards the Schrödinger bridge (3) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t. the reference measure ${\cal P}_{0}$) while the sequence ${\cal P}_{2n+1}$ converges towards the Schrödinger bridge from $\nu_{V}$ to $\lambda_{U}$ (w.r.t. the reference measure ${\cal P}_{1}$).
As shown in Section 6 of [2], under mild integrability assumptions the Schrödinger bridge distribution $\mathbb{P}$ solving (3) can also be computed from any Sinkhorn bridge distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~{}|~{}{\cal P}_{2n}).$ (8)
The Sinkhorn transition operators can also be written in terms of the potential functions $(U_{n},V_{n})$, namely,
${\cal S}_{2n}(x,dy)=\frac{{\cal Q}(x,dy)e^{-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\cal S}_{2n+1}(y,dx)=\frac{{\cal R}(y,dx)~{}e^{-U_{2n+1}(x)}}{{\cal R}(e^{-U_{2n+1}})(y)},$ (9)
while Proposition 6.3 in [2] yields the formulae
$V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$ (10)
for the potentials $(U_{n},V_{n})$, $n\geq 0$, expressed in terms of the distribution flow of the internal states of the Gibbs-loop process.
In addition, for any $n\geq 0$ we also have the commutation formulae
${\cal S}_{2n}\left(\frac{d\nu_{V}}{d\pi_{2n}}\right)=\frac{d\pi_{2n+1}}{d\lambda_{U}}\quad\mbox{and}\quad{\cal S}_{2(n+1)}\left(\frac{d\pi_{2n}}{d\nu_{V}}\right)=\frac{d\lambda_{U}}{d\pi_{2n+1}}.$ (11)
In the same vein, for any $n\geq 1$ we have
${\cal S}_{2n-1}\left(\frac{d\lambda_{U}}{d\pi_{2n-1}}\right)=\frac{d\pi_{2n}}{d\nu_{V}}\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)=\frac{d\nu_{V}}{d\pi_{2n}}.$ (12)
A proof for (11) and (12) is provided in Appendix E on page E.
2.3 Sinkhorn semigroups
For any $n\geq 0$ we have the fixed point equations
$\lambda_{U}{\cal S}^{\circ}_{2n+1}=\lambda_{U}\quad\mbox{\rm and}\quad\nu_{V}{\cal S}^{\circ}_{2(n+1)}=\nu_{V}$ (13)
with the forward-backward Gibbs-type Markov transitions
${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}.$
The above fixed point equations yield, for any $n\geq 1$, the time inhomogeneous Markov evolution equations
$\pi_{2n}=\pi_{2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2n-1}{\cal S}^{\circ}_{2n+1}.$ (14)
For any $0\leq p\leq n$ we have the semigroup equations
$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{2n+1}}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
The Sinkhorn transitions ${\cal S}^{\circ}_{2p,2n}$ and ${\cal S}^{\circ}_{2p+1,2n+1}$ are defined inductively by the time inhomogeneous semigroup equations
${\cal S}^{\circ}_{2p,2n}={\cal S}^{\circ}_{2p,2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2p+1,2n+1}={\cal S}^{\circ}_{2p+1,2n-1}{\cal S}^{\circ}_{2n+1},$
with ${\cal S}^{\circ}_{2n,2n}=I={\cal S}^{\circ}_{2n+1,2n+1}$ the identity integral operator. The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (14) sharing the same invariant measures (13) at each individual step. Using the commutation properties (11) and (12) for any $n\geq p$ we readily check that
$\frac{d\pi_{2n}}{d\nu_{V}}$ $=$ ${\cal S}^{\circ}_{2n}\left(\frac{d\pi_{2(n-1)}}{d\nu_{V}}\right)={\cal S}^{\circ}_{2n,2p}\left(\frac{d\pi_{2p}}{d\nu_{V}}\right)\quad\text{and}$
$\frac{d\pi_{2n+1}}{d\lambda_{U}}$ $=$ ${\cal S}_{2n+1}^{\circ}\left(\frac{d\pi_{2n-1}}{d\lambda_{U}}\right)={\cal S}^{\circ}_{2n+1,2p-1}\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right),$ (15)
with the backward semigroups
${\cal S}^{\circ}_{2n,2p}:={\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1,2p-1}:={\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1}.$
By (11) and (12) we also obtain
$\frac{d\lambda_{U}}{d\pi_{2n+1}}$ $=$ $\left({\cal S}_{2(n+1)}\,{\cal S}^{\circ}_{2n,2p}\,{\cal S}_{2p-1}\right)\left(\frac{d\lambda_{U}}{d\pi_{2p-1}}\right)\quad\text{and}$
$\frac{d\nu_{V}}{d\pi_{2n}}$ $=$ $\left({\cal S}_{2n+1}\,{\cal S}^{\circ}_{2n-1,2p-1}\,{\cal S}_{2(p-1)}\right)\left(\frac{d\nu_{V}}{d\pi_{2(p-1)}}\right).$ (16)
3 Contraction of Markov operators
We recall some definitions of contraction coefficients for Markov operators, both in terms of the Hilbert projective metric and $\Phi$-entropies. These coefficients will be key to the analysis of contraction properties of Sinkhorn semigroups in Section 4.
3.1 Metrics and divergences
3.1.1 Hilbert projective metric
The Hilbert (projective) metric is defined, for any comparable measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$, by the formula
$H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~{}\|\nu_{1}-\nu_{2}\|_{\sf tv},$ (17)
[…]
3.1.2 Integral operator norms
Given some functions $\varphi\in{\cal B}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\cal M}_{\varphi}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\cal M}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}$ ::=:= $\sup\{|\nu(f)|~{}:~{}\|f\|_{\varphi}\leq 1\}=\sup\{|\nu(f)|~{}:~{}\mbox{\rm osc}_{\varphi}(f)\leq 1\}$ (18)
$=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation}\quad|\nu|:=\nu_{+}+\nu_{-}.$
In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\cal B}(\mathbb{X})$, which is given by
$\mbox{\rm osc}_{\varphi}(f):=\sup_{x_{1},x_{2}}\frac{|f(x_{1})-f(x_{2})|}{\varphi(x_{1})+\varphi(x_{2})}\leq\|f\|_{\varphi}.$
For a detailed proof of the equivalent formulations in the latter definition (18), we refer to Proposition 8.2.16 in [38].
Whenever $\varphi\geq 1$, for any $\kappa>0$, we can construct the function $\phi_{\kappa}=1+\kappa\phi$ and obtain the norm equivalence formulae
$\kappa~{}\|\nu\|_{\sf tv}\leq\kappa~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi_{\kappa}}\leq(2+\kappa){\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}.$ (19)
The choice of condition $\varphi\geq 1/2$ in the definition of the $\varphi$-norm given in (18) is imposed only to recover the conventional total variation dual distance ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\mu_{1}-\mu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{1/2}=\|\mu_{1}-\mu_{2}\|_{\sf tv}$ between probability measures $\mu_{1},\mu_{2}\in{\cal M}_{1}(\mathbb{X})={\cal M}_{1,\frac{1}{2}}(\mathbb{X})$, where ${\cal M}_{1,\varphi}(\mathbb{X})$ denotes the subset of probability measures $\nu$ s.t. $\nu(\varphi)<\infty$. Recall that for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
$\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\cal M}_{1}(S)~{}:~{}\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$ (20)
3.1.3 Entropy and divergences
Consider a convex function
$\begin{array}[]{cccc}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\\ \end{array}$
such that for any $\alpha,v,w\in\mathbb{R}_{+}$ we have
$\Phi(\alpha w,\alpha v)=\alpha~{}\Phi(u,v)\quad\mbox{\rm and}\quad\Phi(1,1)=0.$ (21)
The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
$D_{\Phi}(\nu_{1},\nu_{2}):=\int~{}\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~{}\gamma(dx).$
By homogeneity arguments, the above definition does not depend on the choice of the dominating measure $\gamma$. Note that
$\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~{}|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$ (22)
Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [26], i.e., we also have
$D_{\Phi}(\nu_{1},\nu_{2})=\int~{}\phi\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~{}\nu_{2}(dx)\quad\text{where}\quad\phi(u):=\Phi(u,1)$
and we have used the convention $D_{\Phi}(\nu_{1},\nu_{2})=\infty$ when $\nu_{1}\not\ll\nu_{2}$. Therefore, the $\Phi$-entropy generalizes the classical relative entropy (a.k.a. Kullback-Leibler divergence or $I$-divergence). More precisely, choosing the functions
$\Phi(u,v)=\Phi_{1}(u,v):=u\log{(u/v)}\quad\mbox{\rm and}\quad\phi(u)=u\log{u}$
we recover the classical formulae
$D_{\Phi_{1}}(\nu_{1},\nu_{2})=\mbox{\rm Ent}(\nu_{1}~{}|~{}\nu_{2}):=\int~{}\log{\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)}~{}\nu_{1}(dx)=\int~{}\phi_{1}\left(\frac{d\nu_{1}}{d\nu_{2}}(x)\right)~{}\nu_{2}(dx),$
with the convex function with unique minimum $\phi_{1}(1)=0=\partial\phi_{1}(1)$ defined by
$\phi_{1}(u)=u\log{u}-(u-1).$
Other examples of $\phi$-divergences includes $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [26, 37, 63, 65].
3.2 Some contraction coefficients
Let $\delta_{x}$ denote the Dirac delta measure located at $x$. Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$, we denote by $\imath_{x_{1},x_{2}}({\cal K})(y)$ the Radon-Nikodym function
$\imath_{x_{1},x_{2}}({\cal K})(y):=\frac{d\delta_{x_{1}}{\cal K}}{d\delta_{x_{2}}{\cal K}}(y).$ (23)
We also consider the $[0,1]$-valued parameters
$\jmath({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~{}\inf_{y\in\mathbb{Y}}~{}\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
$\hbar({\cal K})$ ::=:= $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (24)
and we note that
$\delta_{x_{1}}{\cal K}(dy_{1})~{}\delta_{x_{2}}{\cal K}(dy_{2})~{}\geq\hbar({\cal K})~{}\delta_{x_{2}}{\cal K}(dy_{1})~{}\delta_{x_{1}}{\cal K}(dy_{2}).$
Integrating the above inequality w.r.t. the variable $y_{2}$ we readily see that
$\delta_{x_{1}}{\cal K}(dy_{1})~{}\geq\hbar({\cal K})~{}\delta_{x_{2}}{\cal K}(dy_{1}),\quad\text{which readily yields}\quad\jmath({\cal K})\geq\hbar({\cal K})\geq\jmath({\cal K})^{2}.$ (25)
3.2.1 Birkhoff contraction coefficient
The Birkhoff contraction coefficient of a Markov transition ${\cal K}$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
[…]
where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B. Combining (25) and (26) we readily obtain the equivalent inequalities
[…]
As remarked in Section 3 of [22], the Hilbert metric dominates all $\phi$-divergences, thus we expect rather strong conditions to ensure the contraction of Sinkhorn transitions w.r.t. the Hilbert metric.
3.2.2 Dobrushin contraction coefficients
[…]
$\chi_{\varphi,\psi}({\cal K})=\sup_{\nu_{1},\nu_{2}\in{\cal M}_{1,\varphi}(S)}\frac{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}{\cal K}-\nu_{2}{\cal K}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu_{1}-\nu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}}$
When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\cal K})$ instead of $\chi_{\varphi,\varphi}({\cal K})$. Given a Markov transition ${\cal K}_{0}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a Markov transition ${\cal K}_{1}(x,dy)$ from $\mathbb{Y}$ into $\mathbb{X}$ we have
$\chi_{\varphi}({\cal K}_{0}{\cal K}_{1})=\chi_{\varphi,\varphi}({\cal K}_{0}{\cal K}_{1})\leq\chi_{\varphi,\psi}({\cal K}_{0})~{}\chi_{\psi,\varphi}({\cal K}_{1}).$ (27)
[…]
with the parameter
$\varepsilon({\cal K}):=1-\sup_{(x_{1},x_{2})\,\in\mathbb{X}^{2}}{\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}}.$ (29)
Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$ using the estimate (25) as well as (20) we readily find that
$\varepsilon({\cal K})\geq\jmath({\cal K})\geq\hbar({\cal K})\quad\mbox{\rm and}\quad\raisebox{2.0pt}{$\chi$}({\cal K})\leq 1-\jmath({\cal K})\leq 1-\hbar({\cal K}).$ (30)
3.2.3 Entropy contraction coefficient
[…]
For any convex function $\Phi$ satisfying (21) we recall that
[…]
The estimate (32) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [21] for finite state spaces and [37] for general measurable spaces). For further details on the Dobrushin contraction coefficient we refer the reader to the article [41], see also [35, 37] and Chapter 4 in [34].
4 Sinkhorn contraction properties
4.1 Some regularity conditions
The $r$-sub-level set of a positive function $\varphi\in{\cal B}(\mathbb{X})$ is denoted as
$C_{\varphi}(r,\mathbb{X}):=\left\{x\in\mathbb{X}~{}:~{}\varphi(x)\leq r\right\}.$
When there is no possible confusion we write $C_{\varphi}(r)=\{\varphi\leq r\}$ instead of $C_{\varphi}(r,\mathbb{X})$. Lyapunov functions and local minorization conditions are often used in the analysis of time inhomogeneous Markov chains (see [6] and references therein). In the context of Gibbs-loop semigroups (14), such conditions are expressed in terms of some Lyapunov-type functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi\wedge\psi\geq 1$, as follows.
Condition (${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$).
There exist some parameters $\epsilon\in]0,1[$ some constant $c>0$ and some integer $n_{0}\geq 1$ such that $\pi_{2n_{0}}(\psi)\vee\pi_{2n_{0}+1}(\varphi)<\infty$ and for any $n\geq n_{0}$ we have the estimates
${\cal S}_{2n}^{\circ}(\psi)\leq\epsilon~{}\psi+c\quad\mbox{and}\quad{\cal S}_{2n+1}^{\circ}(\varphi)\leq\epsilon~{}\varphi+c.$ (33)
There also exists some $r_{0}\geq 1$ and some function
$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}^{\circ}_{2n+1}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}^{\circ}_{2n}\|_{\sf tv}\leq 1-\varepsilon(r).$ (34)
The inequality (34) is satisfied if, and only if, there exists some ${\cal X}\in{\cal M}_{1}(\mathbb{X})$ (respectively, ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$) that may depend on $(n,r,(x_{1},x_{2}))$ (respectively, on $(n,r,(y_{1},y_{2}))$) satisfying, for any $i\in\{1,2\}$, the minorization conditions
${\cal S}^{\circ}_{2n}(y_{i},dy)\geq\varepsilon(r)~{}{\cal Y}(dy)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}(x_{i},dx)\geq\varepsilon(r)~{}{\cal X}(dx).$ (35)
The above assertion is rather well known and it can be checked using (20). Note that the drift condition (33) is trivially met for the unit functions $\varphi=1=\psi$ but in this case for any $r\geq 1$ we have $C_{\varphi}(r)=\mathbb{X}$ and $C_{\psi}(r)=\mathbb{Y}$. In that case, the inequalities in (35) are uniform minorization conditions over $\mathbb{X}$ and $\mathbb{Y}$.
Let ${\cal B}_{\infty}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the subalgebra of locally bounded and uniformly positive functions $\varphi$ that grow at infinity; that is, $\sup_{C}\varphi<\infty$ for any compact set $C\subset\mathbb{X}$, and for any
$r\geq\varphi_{\star}:=\inf_{\mathbb{X}}\varphi>0$ (36)
the $r$-sub-level set $C_{\varphi}(r)$ is a non empty compact subset. We denote by ${\cal B}_{0}(\mathbb{X}):=\{1/\varphi~{}:~{}\varphi\in{\cal B}_{\infty}(\mathbb{X})\}\subset{\cal B}_{b}(\mathbb{X})$ the sub-algebra of (bounded) positive functions, locally lower bounded and that vanish at infinity. When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ the $r$-sub-level sets $C_{\varphi}(r)$ and $C_{\psi}(r)$ are compact and the inequality (35) yields local minorization conditions.
The assumption $\varphi\wedge\psi\geq 1$ is only made to simplify the presentation. Indeed, for any function $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi\in{\cal B}_{\infty}(\mathbb{Y})$ satisfying (33), the rescaled functions $\varphi/\varphi_{\star}\in{\cal B}_{\infty}(\mathbb{X})$ and $\psi/\psi_{\star}\in{\cal B}_{\infty}(\mathbb{Y})$ also satisfy (33).
The drift conditions in (33) ensure that for any $n\geq n_{0}$ we have
${\cal S}^{\circ}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}^{\circ}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\varphi}(\mathbb{X}).$
Whenever Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, it can be easily shown that
$\lambda_{U}(\varphi)\vee\nu_{V}(\psi)<\infty\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))<\infty.$ (37)
These estimates are direct consequences of the drift conditions (33). A detailed proof is provided in Section E on page E.
4.2 Norm contraction theorems
The drift conditions (33) combined with the local minorization conditions (34) provide a way to control the contraction coefficients of Sinkhorn transitions. More precisely, we have the following estimate for the contraction coefficients of the forward-backward Markov transitions ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$.
Theorem 4.1.
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied, then there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that
$\sup_{n\geq n_{0}}\left(\raisebox{2.0pt}{$\chi$}_{\psi_{\kappa}}({\cal S}^{\circ}_{2n})\vee\raisebox{2.0pt}{$\chi$}_{\varphi_{\kappa}}({\cal S}^{\circ}_{2n+1})\right)\leq\rho,\quad\text{where}\quad(\varphi_{\kappa},\psi_{\kappa}):=(1+\kappa\varphi,1+\kappa\psi).$ (38)
Theorem 4.1 is a straightforward consequence of Theorem C.1 in Appendix C.
The contraction estimate (38) ensures the exponential decay of Sinkhorn semigroups. For instance, for any $\mu_{1},\mu_{2}\in{\cal M}_{1,\psi}(\mathbb{Y})$ and $0\leq p\leq n$ we readily obtain
$\|(\mu_{1}-\mu_{2}){\cal S}^{\circ}_{2p,2n}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\mu_{1}-\mu_{2}\|_{\psi_{\kappa}}.$
Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\pi_{0},\nu_{V})$, the above estimate immediately yields an upper bound on the total variation distance, i.e.,
$\|\pi_{2n}-\nu_{V}\|_{\sf tv}\leq\|\pi_{2n}-\nu_{V}\|_{\psi_{\kappa}}\leq\rho^{n-p}\|\pi_{0}-\nu_{V}\|_{\psi_{\kappa}}.$
Choosing $p=0$ and $(\mu_{1},\mu_{2})=(\delta_{y_{1}},\delta_{y_{2}})$ for some $(y_{1},y_{2})\in\mathbb{Y}^{2}$, we can also verify the merging property of Sinkhorn semigroups [30, 70, 71]
$\|{\cal S}^{\circ}_{0,2n}(y_{1},\mbox{\LARGE.})-{\cal S}^{\circ}_{0,2n}(y_{2},\mbox{\LARGE.})\|_{\sf tv}\leq\rho^{n}~{}\left(\psi(y_{1})+\psi(y_{2})\right).$
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is expressed in terms of the forward-backward transition kernels ${\cal S}^{\circ}_{2n}$ and ${\cal S}^{\circ}_{2n+1}$, which may make it hard to test –except in the case of Gaussian and linear models. Several Lyapunov functions in this context are designed in Section 5.2. A stronger (but easier to test) condition can be provided in terms of the Sinkhorn transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$.
Condition (${\cal D}_{\varphi,\psi}({\cal S})$).
There exists some parameter $\epsilon\in]0,1[$, some constant $c>0$ and some integer $n_{0}\geq 0$ such that for any $n\geq n_{0}$ we have the estimates
${\cal S}_{2n}(\psi)\leq\epsilon~{}\varphi+c\quad\mbox{and}\quad{\cal S}_{2n+1}(\varphi)\leq\epsilon~{}\psi+c.$ (39)
There also exists some $r_{0}\geq 1$ and some function
$\begin{array}[]{cccc}\varepsilon:&[r_{0},\infty[&\mapsto&]0,1]\\ &r&\leadsto&\varepsilon(r)\\ \end{array}$
such that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$, $(y_{1},y_{2})\in C_{\psi}(r)^{2}$ and $n\geq n_{0}$ we have the uniform estimates
$\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}\|_{\sf tv}\vee\|(\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (40)
When Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is met, we readily see that
${\cal S}_{2n}\left({\cal B}_{\psi}(\mathbb{Y})\right)\subset{\cal B}_{\varphi}(\mathbb{X})\quad\mbox{\rm and}\quad{\cal S}_{2n+1}\left({\cal B}_{\varphi}(\mathbb{X})\right)\subset{\cal B}_{\psi}(\mathbb{Y})$
for any $n\geq n_{0}$.
Theorem 4.2.
If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$, then (33) and (34) also hold for any $n>n_{0}$. In addition, there exist parameters $\kappa\geq 0$ and $\rho\in[0,1[$ (depending only on $\varphi$, $\psi$ and $\varepsilon$) such that, for any $n\geq n_{0}$,
[…]
A proof of Theorem 4.2 is provided in Appendix C (on page C).
4.3 A Lyapunov stability theorem
Finally, we introduce a more tractable sufficient condition, in terms of the potential functions $(U,V,W)$. It enables us to obtain geometric rates for the convergence of the Sinkhorn iterations, as explicitly shown for the case of linear Gaussian models and finite mixture models in Sections 5.2 and 5.3, respectively. Let us assume that $W$ is locally bounded, $(U,V)\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y}))$ and $e^{-W}\in\mathbb{L}_{\infty}(\lambda\otimes\nu)$. In addition, assume that $U$ and $V$ have compact sub-levels sets with non empty interior and the measures $\lambda$ and $\nu$ have full support, i.e., they are finite on compact sets and strictly positive on non-empty open sets. Let us also denote
$W_{U}(y):=\int~{}\lambda_{U}(dx)~{}W(x,y)\quad\text{and}\quad W^{V}(x):=\int~{}\nu_{V}(dy)~{}W(x,y).$ (41)
Condition (${\cal H}_{\delta}(U,V)$).
There is some $\delta\in]0,1[$ such that ${\cal U}_{\delta}:=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$ and ${\cal V}_{\delta}:=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$. In addition, the functions $W_{U},W^{V}$ are locally bounded and
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$
Theorem 4.3.
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$, where $U_{\star}=\inf_{\mathbb{X}}U(x)$ and $V_{\star}=\inf_{\mathbb{Y}}V(y)$. In addition, there exist some parameter $\rho\in[0,1[$ and some constant $c>0$ such that
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c~{}\rho^{2n}\quad\text{for any $n\geq 1$.}$ (42)
A proof of Theorem 4.3 is provided in Appendix D.3.
Remark 4.4.
Using Proposition D.3 in Appendix D and following exactly the same argument as in the proof of Theorem 4.3, the estimate (42) remains valid when the weaker condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, provided that
${}\left(\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\right)\vee\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right)<\infty.$
Let us discuss some fairly straightforward consequences of the estimate (42). First note that (42) ensures the convergence of the series
$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ ::=:= $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ ::=:= $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (43)
where $U_{0}=U$ and $V_{0}=0$. Applying Lebesgue’s dominated convergence theorem, the integral equations stated in (7) converge as $n\rightarrow\infty$ to a system of integral equations
$\mathbb{U}=U+\log{{\cal Q}(e^{-\mathbb{V}})}\quad\mbox{\rm and}\quad\mathbb{V}=V+\log{{\cal R}(e^{-\mathbb{U}})}.$
More precisely, using the estimates in (42) it is straightforward to prove the following corollary.
Corollary 4.5.
Under the assumptions of Theorem 4.3 for any $n\geq 1$ we have
$\|\mathbb{U}-U_{2n}\|_{\psi}\vee\|\mathbb{V}-V_{2n}\|_{\varphi}\leq\frac{c}{1-\rho^{2}}~{}\rho^{2n}$ (44)
with the constant $c>0$ as in (42). In addition, we have the relative entropy estimates
$\mbox{\rm Ent}\left(\lambda_{U}~{}|~{}\pi_{2n+1}\right)\leq c~{}\lambda_{U}(\varphi)~{}\rho^{2n}~{}\quad\mbox{and}\quad\mbox{\rm Ent}\left(\nu_{V}~{}|~{}\pi_{2n}\right)\leq c~{}\nu_{V}(\psi)~{}\rho^{2n}.$ (45)
With this notation at hand, the bridge distribution (3) has the form
$\lim_{n\to\infty}{\cal P}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~{}\lambda_{\mathbb{U}}(dx)~{}\nu_{\mathbb{V}}(dy).$ (46)
In addition, we have
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right)$
The exponential estimate (42) also allows the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (45).
Corollary 4.6.
Under the assumptions of Theorem 4.3 we obtain
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})$ $=$ $\sum_{p\geq n}\mbox{\rm Ent}\left(\lambda_{U}~{}|~{}\pi_{2p+1}\right)+\sum_{p\geq n}\mbox{\rm Ent}\left(\nu_{V}~{}|~{}\pi_{2p}\right)\quad\text{for any $n\geq 1$}.$
In addition, there exists some $c>0$ such that, for any $n\geq 1$, we have the entropy exponential decay estimates
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})\leq c~{}\rho^{2n}~{}(\lambda_{U}(\varphi)+\nu_{V}(\psi)).$
5 Some illustrations
5.1 Strong mixing models
As underlined in [45, 62] in the context of finite state space models, Sinkhorn iterations of the form (5) preserve the cross product Radon-Nikodym ratio in (24) and, hence, we have
$\hbar({\cal S}_{n})=\hbar({\cal S}_{0})=\hbar({\cal Q}),\quad\forall n\geq 0.$ (47)
This property is crucial in the convergence analysis of Sinkhorn iterations using Hilbert-projective techniques, see for instance [14, 18, 33, 47]. Recall that the Hilbert contraction coefficient of a Markov transition ${\cal Q}$ can be explicitly computed in terms of $\hbar({\cal Q})$ (see Eq. (26)), hence the exponential convergence of Sinkhorn iterations w.r.t. the Hilbert metric takes place if, and only if, $\hbar({\cal Q})>0$. An elementary proof of this property is presented in Proposition B.1 (in Appendix B).
From (25), let us observe that
$\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})\geq\jmath({\cal Q})^{2}.$
Thus, for any given a convex function $\Phi$ satisfying (21), we can use (26) and (32) to show that
$\hbar({\cal Q})>0\quad\text{if and only if}\quad\jmath({\cal Q})>0,$
while
$\jmath({\cal Q})>0\quad\text{implies}\quad\left\{\begin{array}[]{l}\raisebox{2.0pt}{$\chi$}_{H}({\cal S}_{n})=\raisebox{2.0pt}{$\chi$}_{H}({\cal Q})<1,\\ \text{${\cal D}_{1,1}({\cal S}^{\circ})$ is satisfied with $\varepsilon(r)=\hbar({\cal Q})>0$ for any $r\geq 1$, and}\\ \sup_{n\geq 0}\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal S}_{n})\leq\raisebox{2.0pt}{$\chi$}({\cal S}):=\sup_{n\geq 0}\raisebox{2.0pt}{$\chi$}({\cal S}_{n})\leq 1-\hbar({\cal Q})<1.\end{array}\right.$
We can also obtain the sandwich-type contraction inequalities described below.
Theorem 5.1.
For any convex function $\Phi$ satisfying (21) we have the contraction estimates
$\begin{array}[]{rcccl}D_{\Phi}\left(\pi_{2(n+1)},\nu_{V}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~{}D_{\Phi}\left(\lambda_{U},\pi_{2n+1}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~{}D_{\Phi}\left(\pi_{2n},\nu_{V}\right),\\ &&&&\\ D_{\Phi}\left(\nu_{V},\pi_{2(n+1)}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})~{}D_{\Phi}\left(\pi_{2n+1},\lambda_{U}\right)&\leq&\raisebox{2.0pt}{$\chi$}({\cal S})^{2}~{}D_{\Phi}\left(\nu_{V},\pi_{2n}\right).\end{array}$ (48)
[…]
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|\leq c~{}\raisebox{2.0pt}{$\chi$}({\cal S})^{2n}\quad\text{for any $n\geq 0$.}$ (49)
Proof.
The estimates (48) are a direct consequence of (32) combined with the Markov transport formulae
$(\pi_{2(n+1)},\nu_{V})$ $=$ $(\lambda_{U}{\cal S}_{2(n+1)},\pi_{2n+1}{\cal S}_{2(n+1)}),\quad\text{and}$
$(\lambda_{U},\pi_{2n+1})$ $=$ $(\pi_{2n}{\cal S}_{2n+1},\nu_{V}{\cal S}_{2n+1}),$
which are themselves easily obtained from expressions (5) and (13). The proof of (49) follows exactly the same arguments as the proof of the inequality (42) in Theorem 4.3, hence it is skipped.
The following result can be obtained from the series expansion given by Corollary 4.6.
Corollary 5.2.
Under the assumptions of Theorem 5.1, we have the entropy exponential decay estimates
$\mbox{\rm Ent}(\mathbb{P}~{}|~{}{\cal P}_{2n})\leq c~{}\raisebox{2.0pt}{$\chi$}({\cal S})^{2n},\quad\text{for any $n\geq 0$}.$
Several examples satisfying condition $\jmath({\cal Q})>0$ are discussed in Section 2.4 in [35]. For instance, consider a parameter $\delta>0$ and drift function $x\mapsto b(x)$ with bounded oscillations from $\mathbb{X}=\mathbb{R}^{d}$ into itself. The choice ${\cal Q}(x,dy)~{}\propto~{}e^{-\delta\,\|y-b(x)\|}~{}dx$ yields
$\jmath({\cal Q})\geq\exp{\left(-\delta\,\Delta(b)\right)},\quad\text{with}\quad\Delta(b)=\sup_{(x_{1},x_{2})\in\mathbb{X}^{2}}{\|b(x_{1})-b(x_{2})\|}.$
In continuous time settings, condition $\jmath({\cal Q})>0$ is also satisfied for Markov transitions of elliptic diffusions on compact manifolds $\mathbb{X}$ –see for instance the pioneering work of Aronson [7], Nash [64] and Varopoulos [78] on Gaussian estimates for heat kernels on manifolds. Nevertheless, the above condition is not met for linear Gaussian transitions.
5.2 Linear-Gaussian models
[…]
Let ${\cal G}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\cal S}^{+}_{d}\subset{\cal G}l_{d}$ the subset of positive definite matrices. We sometimes use the Löwner partial ordering notation $v_{1}\geq v_{2}$ to mean that a symmetric matrix $v_{1}-v_{2}$ is positive semi-definite (equivalently, $v_{2}-v_{1}$ is negative semi-definite), and $v_{1}>v_{2}$ when $v_{1}-v_{2}$ is positive definite (equivalently, $v_{2}-v_{1}$ is negative definite). Given $v\in{\cal S}_{d}^{+}$ we denote by $v^{1/2}$ the principal (unique) symmetric square root.
Denote by $\nu_{m,\sigma}$ the Gaussian distribution on $\mathbb{R}^{d}$ with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\sigma\in{\cal S}^{+}_{d}$. In addition, let $g_{\sigma}$ denote the probability density function of the distribution $\nu_{0,\sigma}$, with covariance matrix $\sigma\in{\cal S}^{+}_{d}$. Consider the target probability measures $(\lambda_{U},\nu_{V})=(\nu_{m,\sigma},\nu_{\overline{m},\overline{\sigma}})$ for some given parameters
$(m,\overline{m})\in(\mathbb{R}^{d}\times\mathbb{R}^{d})\quad\text{and}\quad(\sigma,\overline{\sigma})\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d}).$
Equivalently, the log-density functions $(U,V)$ are defined by
$U(x)=-\log{g_{\sigma}(x-m)}\quad\mbox{\rm and}\quad V(y):=-\log{g_{\overline{\sigma}}(y-\overline{m})}.$ (50)
The linear Gaussian transition ${\cal Q}$ is defined by (2), with the potential
$W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$ (51)
for some given $(\alpha,\beta,\tau)\in\left(\mathbb{R}^{d}\times{\cal G}l_{d}\times{\cal S}^{+}_{d}\right)$. In this context, we have $\jmath({\cal Q})=0$ and Hilbert projective techniques cannot be applied. However, a refined exponential stability analysis of the Sinkhorn algorithm for Gaussian models has been developed in [2]. For instance, following Remark 5.4 in [2], there exists some $n_{0}\geq 1$ such that for any $n\geq n_{0}$ we have the estimate
$\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~{}\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$ (52)
The matrix $\Delta_{\varpi}$ in expression (52) above is defined as
$\Delta_{\varpi}:=\frac{\varpi}{2}+\left(\varpi+\left(\frac{\varpi}{2}\right)^{2}\right)^{1/2}\geq\varpi+(I+\varpi^{-1})^{-1}$ (53)
with
$\varpi:=(\gamma\,\gamma^{\prime})^{-1}\quad\mbox{\rm and}\quad\gamma:=\overline{\sigma}^{1/2}~{}\tau^{-1}\beta~{}\sigma^{1/2}.$
The drift-minorization techniques developed in this paper also enable us to obtain exponential convergence on weighted Banach spaces.
Proposition 5.3.
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied with the Lyapunov functions
$\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$ (54)
A proof is provided in Appendix A (on page A).
The choice of the Lyapunov functions is not unique. For instance, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
$\varphi(x)$ ::=:= $\exp{\left(\delta\,((x-m)^{\prime}\sigma^{-1}(x-m))^{1/2}\right)}\quad\text{and}$
$\psi(y)$ ::=:= $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (55)
There also exists some $\delta_{0}>0$ such that ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
$\varphi(x)$ ::=:= $\exp{\left(\delta\,(x-m)^{\prime}\sigma^{-1}(x-m)\right)}\quad\text{and}$
$\psi(y)$ ::=:= $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (56)
A detailed proof of the above assertions is provided in Appendix A on page A.
We end this section with an equivalent formulation of condition ${\cal H}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (50) and (51).
Proposition 5.4.
Condition ${\cal H}_{\delta}(U,V)$ is satisfied for the Lyapunov functions in (56) if, and only if,
$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~{}\sigma^{-1}\quad\text{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~{}\overline{\sigma}^{-1}.$ (57)
A proof is provided in Appendix E (on page E).
5.3 Finite mixture models
Consider some (fully supported) discrete probability measures $\xi_{0}$ and $\xi_{1}$ on some some finite states $E_{0}$ and $E_{1}$ and some collection of probability measures $(\lambda_{U_{i}})_{i\in E_{0}}$ and $(\nu_{V_{j}})_{j\in E_{1}}$ for some continuous potential functions $U_{i}:x\in\mathbb{X}\mapsto U(x)\in\mathbb{R}$ and $V_{j}:y\in\mathbb{X}\mapsto V(y)\in\mathbb{R}$ indexed by $i\in E_{0}$ and $j\in E_{1}$. The mixture distributions $\lambda_{U}$ and $\nu_{V}$ associated with these potential functions take the form
$U=-\log{\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}}\quad\mbox{\rm and}\quad V=-\log{\sum_{j\in E_{1}}\xi_{1}(j)~{}e^{-V_{j}}}.$ (58)
Proposition 5.5.
Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1[$ if conditions ${\cal H}_{\delta}(U_{i},V_{j})$ are met for the same $\delta\in]0,1[$ and almost every $(i,j)\in(E_{0}\times E_{1})$.
A proof is provided in Appendix E (on page E).
Let us now consider the mixture model (58) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (51) and the log-densities
$-U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (59)
where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters. Combining Proposition 5.5 with Proposition 5.4 we readily obtain the following result.
Proposition 5.6.
The Gaussian mixture model defined in (58) with the Gaussian log-densities of (59) defined on $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$ satisfies Condition ${\cal H}_{\delta}(U,V)$ for some $\delta\in]0,1[$ provided that
$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~{}\sigma_{i}^{-1}\quad\mbox{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~{}\overline{\sigma}_{j}^{-1}$ (60)
for almost every $(i,j)\in(E_{0}\times E_{1})$.
Let $(X_{i})_{1\leq i\leq N}$ be $N$ independent and identically distributed samples from some (arbitrary) data distribution $\eta(dx)$ on $\mathbb{X}=\mathbb{R}^{d}$. The Gaussian-kernel estimation model defined by
$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~{}dx.$ (61)
is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$). The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing. Note that (61) coincides with (58) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$. For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [67, 69] as well as the books [16, 73, 79].
The Gaussian mixture distribution (61) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [31] and references therein. Choosing $(\alpha,\beta)=(0,I)$ and $\tau=tI$; for some $t>0$ in (51), the Markov transition ${\cal Q}(x,dy)$ coincides with the the transition semigroup of the $d$-dimensional Brownian motion, namely
${\cal Q}(x,dy)=(2\pi t)^{-d/2}~{}\exp{\left(-\frac{1}{2t}~{}\|y-x\|^{2}\right)}~{}dy.$
In this context, the l.h.s. condition in (60) applied to (61) becomes
$\frac{1}{t}~{}I<(\delta\wedge(1-\delta))~{}\sigma^{-1}.$ (62)
[…]
$e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~{}\sigma^{-1}.$ (63)
Both transitions gradually transform the data distribution $\lambda_{U}$ into a simpler distribution $\lambda_{U}{\cal Q}$ by injecting noise into data. Also note that (62) and (63) are met when the time horizon is chosen sufficiently large.
The choice of the prior distribution $\nu_{V}$ is rather flexible. For instance, we can choose a Gaussian distribution, say $\nu_{0,\overline{\sigma}}$, satisfying the r.h.s. condition in (57). Starting from such Gaussian random states, the backward Sinkhorn transitions (indexed by odd indices) eventually generates samples back into the initial distribution by slowly removing the noise distortion. Conditions (62) and (63) reflect the fact that the forward noise injection as well as the reverse denoising procedure in diffusion-based generative models often requires very large time horizons.
Acknowledgements
This work is supported by the Innovation and Talent Base for Digital Technology and Finance (B21038) and ‘the Fundamental Research Funds for the Central Universities’, Zhongnan University of Economics and Law (2722023EJ002).
J. Miguez acknowledges the financial support of the Office of Naval Research (award no. N00014-22-1-2647) and MICIU/AEI/10.13039/501100011033/FEDER UE (grant no. PID2021-125159NB-I00 TYCHE).
Appendix A The Gaussian Sinkhorn algorithm
The Frobenius matrix norm of a given matrix $v\in\mathbb{R}^{d_{1}\times d_{2}}$ is defined by $\left\|v\right\|_{F}^{2}=\mbox{\rm Tr}(v^{\prime}v)$, with the trace operator $\mbox{\rm Tr}(\cdot)$ and $v^{\prime}$ the transpose of the matrix $v$. Note that, for $x\in\mathbb{R}^{d\times 1}=\mathbb{R}^{d}$, the Frobenius norm $\|x\|_{F}=\sqrt{x^{\prime}x}$ coincides with the Euclidean norm. The spectral norm is defined by $\|v\|_{2}=\sqrt{\lambda_{\sf max}(v^{\prime}v)}$.
Riccati difference equations
Let us associate with some $\varpi\in{\cal S}^{+}_{d}$ the increasing map $\mbox{\rm Ricc}_{\varpi}$ from ${\cal S}^{0}_{d}$ into ${\cal S}^{+}_{d}$ defined by
$\begin{array}[]{lccl}\mbox{\rm Ricc}_{\varpi}:&{\cal S}^{0}_{d}&\mapsto&{\cal S}^{+}_{d}\\ &v&\leadsto&\mbox{\rm Ricc}_{\varpi}(v):=(I+(\varpi+v)^{-1})^{-1}\\ \end{array}.$ (64)
If we define the matrices
$\varpi^{-1}:=\gamma\,\gamma^{\prime}\quad\mbox{\rm and}\quad\overline{\varpi}^{-1}:=\gamma^{\prime}\gamma\in{\cal S}^{+}_{d},\quad\mbox{\rm with}\quad\gamma:=\overline{\sigma}^{1/2}~{}\tau^{-1}\beta~{}\sigma^{1/2},$
and $\upsilon_{0}:=\overline{\sigma}^{-1/2}\tau~{}\overline{\sigma}^{-1/2}$, then it can be shown (see Theorem 4.3 in [2]) that the recursions
$\upsilon_{2n+1}^{-1}=I+\gamma^{\prime}~{}\upsilon_{2n}~{}\gamma\quad\mbox{and}\quad\upsilon_{2(n+1)}^{-1}=I+\gamma~{}\upsilon_{2n+1}~{}\gamma^{\prime},$ (65)
satisfy the Riccati difference equations
$\upsilon_{2(n+1)}=\mbox{\rm Ricc}_{\varpi}\left(\upsilon_{2n}\right)\quad\mbox{and}\quad\upsilon_{2n+1}=\mbox{\rm Ricc}_{\overline{\varpi}}\left(\upsilon_{2n-1}\right),\quad\text{for $n\geq 0$}.$ (66)
Let us also define
$\tau_{2n}:=\overline{\sigma}^{1/2}~{}\upsilon_{2n}~{}\overline{\sigma}^{1/2}\quad\mbox{and}\quad\tau_{2n+1}:=\sigma^{1/2}~{}\upsilon_{2n+1}~{}\sigma^{1/2},$
as well as the gain matrices
$\beta_{2n}:=\tau_{2n}~{}\tau^{-1}\beta\quad\mbox{and}\quad\beta_{2n+1}:=\tau_{2n+1}~{}\beta^{\prime}\tau^{-1}.$
The recursions for $\tau_{n}$ and $\beta_{n}$ are initialized with $(\beta_{0},\tau_{0})=(\beta,\tau)$, where $(\beta,\tau)$ are the parameters of the reference Gaussian measure.
Sinkhorn transitions
We also denote by $G$ a $d$-dimensional centred Gaussian random variable (r.v.) with unit covariance. The Sinkhorn transitions in (5) can be expressed as
${\cal S}_{2n}(x,dy)=\mathbb{P}\left({\cal Z}_{2n}(x)\in dy\right)\quad\mbox{\rm and}\quad{\cal S}_{2n+1}(y,dx)=\mathbb{P}\left({\cal Z}_{2n+1}(y)\in dx\right)$
with random maps defined, for any $n\geq 0$, by
${\cal Z}_{2n}(x)$ $=$ $m_{2n}+\beta_{2n}(x-m)+\tau_{2n}^{1/2}~{}G\quad\text{and}$
${\cal Z}_{2n+1}(y)$ $=$ $m_{2n+1}+\beta_{2n+1}(y-\overline{m})+\tau_{2n+1}^{1/2}~{}G.$
The parameters $m_{n}$ in the latter expressions are defined, for any $n\geq 0$, by the recursion
$m_{2n+1}$ $=$ $m+\beta_{2n+1}(\overline{m}-m_{2n}),$
$m_{2(n+1)}$ $=$ $\overline{m}+\beta_{2(n+1)}~{}(m-m_{2n+1})$
with initial condition $m_{0}=\alpha_{0}+\beta_{0}m$, where, again, $(\alpha_{0},\beta_{0})=(\alpha,\beta)$ are parameters of the reference Gaussian measure. Note that $\pi_{n}=\nu_{m_{n},\sigma_{n}}$ with
$\sigma_{2n}:=\beta_{2n}~{}\sigma~{}\beta_{2n}^{\prime}+\tau_{2n}\quad\mbox{and}\quad\sigma_{2n+1}:=\beta_{2n+1}~{}\overline{\sigma}~{}\beta_{2n+1}^{\prime}+\tau_{2n+1}.$
Gibbs-loop process
The transitions of the Gibbs-loop process introduced in Section 2.3 can also be rewritten as
${\cal S}^{\circ}_{2n}(y_{1},dy_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n}(y_{1})\in dy_{2}\right)\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2n+1}(x_{1},dx_{2}):=\mathbb{P}\left({\cal Z}^{\circ}_{2n+1}(x_{1})\in dx_{2}\right)$
with the random maps
${\cal Z}^{\circ}_{2n}(y)$ ::=:= $\overline{m}+\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G$
${\cal Z}^{\circ}_{2n+1}(x)$ ::=:= $m+\beta_{2n+1}^{\circ}(x-m)+\left(\tau_{2n+1}^{\circ}\right)^{1/2}~{}G$
defined in terms of the parameters
$\begin{array}[]{rclcrcl}\beta_{2n+1}^{\circ}&:=&\beta_{2n+1}~{}\beta_{2n},&&\tau_{2n+1}^{\circ}&:=&\tau_{2n+1}+\beta_{2n+1}~{}\tau_{2n}~{}\beta_{2n+1}^{\prime},\\ &&&&&&\\ \beta^{\circ}_{2n}&:=&\beta_{2n}~{}\beta_{2n-1},&&\text{and}~{}~{}\tau^{\circ}_{2n}&:=&\tau_{2n}+\beta_{2n}~{}\tau_{2n-1}~{}\beta_{2n}^{\prime}.\end{array}$
By Lemma 4.9 in [2], we have
$0\leq\overline{\sigma}^{-1/2}~{}\beta^{\circ}_{2(n+1)}~{}\overline{\sigma}^{1/2}=I-\upsilon_{2(n+1)}\leq(I+\varpi)^{-1}$
as well as
$0\leq\sigma^{-1/2}~{}\beta_{2n+1}^{\circ}~{}\sigma^{1/2}=I-\upsilon_{2n+1}\leq(I+\overline{\varpi})^{-1}.$
Moreover, Corollary 4.4 in [2] yields the uniform estimates
$\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}\leq\sigma,~{}~{}\forall n\geq 1.$
Lyapunov functions
From the fixed point equations (13) we have
$\overline{\sigma}=\beta^{\circ}_{2n}\overline{\sigma}\left(\beta^{\circ}_{2n}\right)^{\prime}+\tau^{\circ}_{2n}\geq\tau^{\circ}_{2n}\quad\mbox{and}\quad\sigma=\beta^{\circ}_{2n+1}\sigma\left(\beta^{\circ}_{2n+1}\right)^{\prime}+\tau^{\circ}_{2n+1}\geq\tau^{\circ}_{2n+1},$ (67)
which, in turn, imply that
$\overline{\sigma}_{\varpi}:=\overline{\sigma}^{1/2}(I+\varpi^{-1})^{-1}\overline{\sigma}^{1/2}\leq\tau_{2n}^{\circ}\leq\overline{\sigma}\quad\mbox{and}\quad\sigma_{\overline{\varpi}}:=\sigma^{1/2}(I+\overline{\varpi}^{-1})^{-1}\sigma^{1/2}\leq\tau_{2n+1}^{\circ}\leq\sigma.$ (68)
Also note that if $\overline{\sigma}^{-1}\leq(\tau_{2n}^{\circ})^{-1}$ then $(\tau_{2n}^{\circ})^{1/2}\overline{\sigma}^{-1}(\tau_{2n}^{\circ})^{1/2}\leq I$. Moreover, we also readily see that
$\begin{array}[]{l}\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\\ \\ =(y-\overline{m})^{\prime}\left(\beta^{\circ}_{2n}\right)^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)\\ \\ =(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)^{\prime}\left(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}\overline{\sigma}^{1/2}\right)\right]\overline{\sigma}^{-1/2}(y-\overline{m})+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right),\end{array}$
which implies
$\mathbb{E}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]$ $=$ $(y-\overline{m})^{\prime}\overline{\sigma}^{-1/2}\left[I-\upsilon_{2n}\right]^{2}\overline{\sigma}^{-1/2}(y-\overline{m})$ (69)
$+\mbox{\rm Tr}\left(\overline{\sigma}^{-1/2}\tau_{2n}^{\circ}\overline{\sigma}^{-1/2}\right)$
$\leq$ $(1+\lambda_{\sf min}(\varpi))^{-2}~{}(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})+d.$
We are now in position to prove Proposition 5.3.
Proof of Proposition 5.3
Consider the function
$\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$
The estimate in (69) ensures that
${\cal S}_{2n}^{\circ}(\psi)(y)=\mathbb{E}\left[\psi({\cal Z}^{\circ}_{2n}(y))\right]\leq(1+\lambda_{\sf min}(\varpi))^{-2}~{}\psi(y)+\left(\left(1-(1+\lambda_{\sf min}(\varpi))^{-2}\right)+d\right).$
Next, let us observe that
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})=\frac{1}{\sqrt{\mbox{\rm det}(2\pi\tau_{2n}^{\circ})}}\\ \\ \displaystyle\times~{}\exp{\left(-\frac{1}{2}~{}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\right)}~{}dy_{2}\end{array}$
and
$\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left(y_{2}-\overline{m}\right)+2\left(\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left(\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left((y_{2}-\overline{m})\right)+2~{}(1+\lambda_{\sf min}(\varpi^{-1}))~{}\left(y_{1}-\overline{m}\right)^{\prime}(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\left(y_{1}-\overline{m}\right),\end{array}$
where the second inequality follows from (68). Moreover, arguing again as above we arrive at
$\begin{array}[]{l}\displaystyle\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)^{\prime}\left(\tau^{\circ}_{2n}\right)^{-1}\left((y_{2}-\overline{m})-\beta^{\circ}_{2n}(y_{1}-\overline{m}))\right)\\ \\ \displaystyle\leq 2\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)+2~{}(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~{}\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\end{array}$ (70)
The inequality (70) above together with (67) readily yields
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\frac{1}{\sqrt{\mbox{\rm det}(2\pi\overline{\sigma})}}{}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~{}\left(y_{1}-\overline{m}\right)^{\prime}\overline{\sigma}^{-1}\left(y_{1}-\overline{m}\right)\right)}\\ \\ \displaystyle\times~{}\exp{\left(-\left(y_{2}-\overline{m}\right)^{\prime}\overline{\sigma}_{\varpi}^{-1}\left(y_{2}-\overline{m}\right)\right)}~{}dy_{2}.\end{array}$
Thus for any $y_{1}$ such that
$\psi(y_{1})=1+(y_{1}-\overline{m})^{\prime}\overline{\sigma}^{-1}(y_{1}-\overline{m})\leq r$
we have
$\begin{array}[]{l}\displaystyle{\cal S}^{\circ}_{2n}(y_{1},dy_{2})\geq\varepsilon(r)~{}\nu_{\overline{m},\overline{\sigma}_{\varpi}/2}(dy_{2})\end{array}$
where
$\varepsilon(r)=\frac{1}{\sqrt{\mbox{\rm det}(2\overline{\sigma}\overline{\sigma}_{\varpi}^{-1})}}~{}\\ \\ \displaystyle\times\exp{\left(-(1+\lambda_{\sf min}(\varpi^{-1}))(1+\lambda_{\sf min}(\varpi))^{-2}~{}(r-1)\right)}.$
This shows that the function $\psi$, and similarly the function $\varphi$ defined in (54) satisfy the drift-minorization conditions stated in (33) and (34).
Lyapunov functions in (55) and (56)
Using the triangle inequality for Mahalanobis distances we check that
$\begin{array}[]{l}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\\ \\ \leq\left[(\beta^{\circ}_{2n}(y-\overline{m}))^{\prime}\overline{\sigma}^{-1}(\beta^{\circ}_{2n}(y-\overline{m}))\right]^{1/2}+\left[(\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G)^{\prime}\overline{\sigma}^{-1}(\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G)\right]^{1/2}\\ \\ \leq(1+\lambda_{\sf min}(\varpi))^{-1}~{}((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2},\end{array}$
which yields the almost sure estimate
$\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]^{1/2}\leq(1+\lambda_{\sf min}(\varpi))^{-1}~{}((y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}+\left[G^{\prime}G\right]^{1/2}.$
This implies that the function $\psi$, and similarly the function $\varphi$ defined in (55) satisfy the drift-minorization conditions stated in (33) and (34).
The analysis of the Lyapunov functions (56) is based on the following technical lemma.
Lemma A.1.
For any matrices $\sigma>\tau$ and any $z\in\mathbb{R}^{d}$ and $0\leq\delta\leq 1$ we have
$\mathbb{E}\left[\exp{\left(\frac{\delta}{2}(z+\tau^{1/2}G)^{\prime}\sigma^{-1}(z+\tau^{1/2}G)\right)}\right]=\displaystyle\sqrt{\frac{\mbox{\rm det}(\sigma)}{\mbox{\rm det}(\sigma-\delta\tau)}}~{}\exp{\left(\frac{\delta}{2}z^{\prime}(\sigma-\delta\tau)^{-1}z\right)}.$
Proof.
For any $g\in\mathbb{R}^{d}$ we have
$\frac{\delta}{2}~{}(z+g)^{\prime}\sigma^{-1}(z+g)-\frac{1}{2}~{}g^{\prime}\tau^{-1}g=\frac{\delta}{2}~{}z^{\prime}\sigma^{-1}z+\frac{\delta^{2}}{2}~{}z^{\prime}\sigma^{-1}\raisebox{2.0pt}{$\chi$}\sigma^{-1}z-\frac{1}{2}~{}(g-\delta\raisebox{2.0pt}{$\chi$}\sigma^{-1}z)^{\prime}\raisebox{2.0pt}{$\chi$}^{-1}(g-\delta\raisebox{2.0pt}{$\chi$}\sigma^{-1}z),$
with
$0<\raisebox{2.0pt}{$\chi$}:=(\tau^{-1}-\delta\sigma^{-1})^{-1}=-\frac{1}{\delta}~{}\sigma+\frac{1}{\delta}\sigma\tau^{-1}(\tau^{-1}-\delta\sigma^{-1})^{-1}.$
We note that
$\sigma^{-1}+\delta\sigma^{-1}\raisebox{2.0pt}{$\chi$}\sigma^{-1}$ $=$ $\sigma^{-1}+\delta\sigma^{-1}\left(-\frac{1}{\delta}~{}\sigma+\frac{1}{\delta}\sigma\tau^{-1}(\tau^{-1}-\delta\sigma^{-1})^{-1}\right)\sigma^{-1}$
$=$ $\tau^{-1}(\tau^{-1}-\delta\sigma^{-1})^{-1}\sigma^{-1}=(\sigma-\delta\tau)^{-1}$
and finally observe that
$\frac{1}{\mbox{\rm det}(\tau\raisebox{2.0pt}{$\chi$}^{-1})}=\frac{\mbox{\rm det}(\sigma)}{\mbox{\rm det}(\sigma-\delta\tau)}.$
Using Lemma A.1 we check that
$\begin{array}[]{l}\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[({\cal Z}^{\circ}_{2n}(y)-\overline{m})^{\prime}\overline{\sigma}^{-1}({\cal Z}^{\circ}_{2n}(y)-\overline{m})\right]\right)}\right]\\ \\ =\displaystyle\mathbb{E}\left[\exp{\left(\frac{\delta}{2}\left[\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G\right)^{\prime}\overline{\sigma}^{-1}\left(\beta^{\circ}_{2n}(y-\overline{m})+\left(\tau^{\circ}_{2n}\right)^{1/2}~{}G\right)\right]\right)}\right]\\ \\ =\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~{}\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2})}}~{}\exp{\left(\frac{\delta}{2}~{}(y-\overline{m})^{\prime}~{}(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~{}\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}~{}(y-\overline{m})\right)}.\end{array}$
On the other hand, we have
$\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\leq I$
and arguing as above we conclude that
$(\beta^{\circ}_{2n})^{\prime}\left(\overline{\sigma}-\delta~{}\tau^{\circ}_{2n}\right)^{-1}\beta^{\circ}_{2n}$ $=$ $(\overline{\sigma}^{-1/2}\beta^{\circ}_{2n})^{\prime}\left(I-\delta~{}\overline{\sigma}^{-1/2}\tau^{\circ}_{2n}\overline{\sigma}^{-1/2}\right)^{-1}\overline{\sigma}^{-1/2}\beta^{\circ}_{2n}$ (71)
$\leq$ $\frac{1}{1-\delta}~{}(\beta^{\circ}_{2n})^{\prime}\overline{\sigma}^{-1}\beta^{\circ}_{2n}\leq\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}~{}\overline{\sigma}^{-1}.$ (72)
If we now consider the function
$\psi(y):=\exp{\left(\frac{\delta}{2}(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}$
then we have
$\begin{array}[]{l}\displaystyle\frac{1}{\psi(y)}~{}\mathbb{E}\left(\psi\left({\cal Z}^{\circ}_{2n}(y)\right)\right)\\ \\ \leq\displaystyle\frac{1}{\sqrt{\mbox{\rm det}(I-\delta~{}I)}}~{}\exp{\left(-\frac{\delta}{2}~{}\left(1-\frac{1}{1-\delta}(1+\lambda_{\sf min}(\varpi))^{-2}\right)~{}(y-\overline{m})^{\prime}~{}\overline{\sigma}^{-1}~{}(y-\overline{m})\right)}.\end{array}$
Choosing
$0<\delta<1-\frac{1}{(1+\lambda_{\sf min}(\varpi))^{2}}$
we check that
$\frac{1}{(1-\delta)(1+\lambda_{\sf min}(\varpi))^{2}}<1\quad\text{and}\quad\lim_{|y|\to\infty}\frac{1}{\psi(y)}~{}{\cal S}^{\circ}_{2n}(\psi)(y)=0.$
This implies that the function $\psi$, and similarly the function $\varphi$, defined in (56) satisfy the drift-minorization conditions stated in (33) and (34).
Appendix B Hilbert metric contractions
We recall the Hilbert contraction coefficient formula ([9, 11, 12, 50], see also [8, 22, 59])
[…]
Also note that
$H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})=\sup_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)$
with the Radon-Nikodym function $\imath_{x_{1},x_{2}}({\cal K})(y)$ defined in (23). We also have
$\sup_{(x_{1},x_{2})\in S^{2}}H(\delta_{x_{1}}{\cal K},\delta_{x_{2}}{\cal K})=1/\hbar({\cal K})$
with the parameter $\hbar({\cal K})$ defined in (24). Using the fact that
$\tanh\left(\frac{1}{4}~{}\log{u}\right)=\frac{u^{1/4}-u^{-1/4}}{u^{1/4}-u^{-1/4}}=\frac{u^{1/2}-1}{u^{1/2}+1}=\frac{1-u^{-1/2}}{1+u^{-1/2}}$
we conclude that
[…]
Proposition B.1.
Consider the collection of Sinkhorn transitions ${\cal S}_{n}$ defined in (4). For any $n\geq 0$,
$\hbar({\cal S}_{n})=\hbar({\cal Q})\quad\mbox{and}\quad\raisebox{2.0pt}{$\chi$}_{H}({\cal S}_{n})=\frac{1-\sqrt{\hbar({\cal Q})}}{1+\sqrt{\hbar({\cal Q})}}.$
Proof.
For any $n\geq 0$, we have
$\imath_{y_{1},y_{2}}({\cal S}_{2n+1})(x_{1})~{}\imath_{y_{2},y_{1}}({\cal S}_{2n+1})(x_{2})$ $=$ $\frac{(d\delta_{x_{1}}{\cal S}_{2n}/d\pi_{2n})(y_{1})}{(d\delta_{x_{1}}{\cal S}_{2n}/d\pi_{2n})(y_{2})}\frac{(d\delta_{x_{2}}{\cal S}_{2n}/d\pi_{2n})(y_{2})}{(d\delta_{x_{2}}{\cal S}_{2n}/d\pi_{2n})(y_{1})}$
$=$ $\imath_{x_{2},x_{1}}({\cal S}_{2n})(y_{2})~{}\imath_{x_{1},x_{2}}({\cal S}_{2n})(y_{1})$
which, together with (24), yields
$\hbar({\cal S}_{2n+1})=\hbar({\cal S}_{2n}).$
In the same vein, for any $n\geq 1$ we have
$\imath_{x_{1},x_{2}}({\cal S}_{2n})(y_{1})~{}\imath_{x_{2},x_{1}}({\cal S}_{2n})(y_{2})$ $=$ $\frac{(d\delta_{y_{1}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{1})}{(d\delta_{y_{1}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{2})}~{}\frac{(d\delta_{y_{2}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{2})}{(d\delta_{y_{2}}{\cal S}_{2n-1}/d\pi_{2n-1})(x_{1})}$
$=$ $\imath_{y_{2},y_{1}}({\cal S}_{2n-1})(x_{2})~{}\imath_{y_{1},y_{2}}({\cal S}_{2n-1})(x_{1}),$
hence
$\hbar({\cal S}_{2n})=\hbar({\cal S}_{2n-1}).$
Recalling that ${\cal S}_{0}={\cal Q}$, this ends the proof of the proposition.
Appendix C Weighted norm contractions
We associate with some $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ the following condition.
Condition (${\cal L}_{\varphi,\psi}({\cal K})$).
There exists some parameters $\epsilon\in]0,1[$ and $c<\infty$ such that
${\cal K}(\psi)\leq\epsilon~{}\varphi+c.$ (75)
In addition, there exist some constant $r_{0}\geq 1$ and some function
$\begin{array}[]{ccccc}\varepsilon&:&[r_{0},\infty[&\mapsto&]0,1]\\ &&r&\leadsto&\varepsilon(r)\\ \end{array}$
such that, for any $r\geq r_{0}$ and $(x_{1},x_{2})\in C_{\varphi}(r)^{2}$,
$\chi_{\varphi}(r,{\cal K}):=\sup_{(x_{1},x_{2})\in C_{\varphi}(r)^{2}}\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}<1-\varepsilon(r).$ (76)
By (20), condition (76) is equivalent to the existence of some ${\cal Y}\in{\cal M}_{1}(\mathbb{Y})$ (that may depends on $(r,x_{1},x_{2})$) satisfying for $i\in\{1,2\}$ the local minorization condition
${\cal K}(x_{i},dy)\geq\varepsilon(r)~{}{\cal Y}(dy).$ (77)
For instance, assume that $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $x_{1},x_{2}\in\mathbb{X}$ and set
$\jmath_{\varphi,\psi}(r,{\cal K}):=\inf_{(x_{1},x_{2})\in C_{\varphi}(r)}~{}\inf_{y\in C_{\psi}(r)}\imath_{x_{1},x_{2}}({\cal K})(y).$ (78)
In this notation, we see that if $\jmath_{\varphi,\psi}(r,{\cal K})\geq\varepsilon(r)$ then (77) holds.
Note that the strength of conditions (75), (76) depends on the strength of the functions $(\varphi,\psi)$. When the functions are bounded, the geometric drift condition (75) is trivially met, but in this case (77) is a uniform contraction condition on the space $\mathbb{X}$.
When $\varphi\in{\cal B}_{\infty}(\mathbb{X})$ the $r$-sub-level sets $C_{\varphi}(r)$ are compact condition (77) is met as soon as
${\cal K}(x,dy)\geq q(x,y)~{}\nu(dy)$ (79)
for some Radon positive measure $\nu$ on $\mathbb{Y}$ and some density function $q$, satisfying the local minorization condition
$\inf_{C_{\varphi}(r)\times C_{\psi}(r)}q>0\quad\mbox{\rm and}\quad 0<\nu(C_{\psi}(r))<\infty$ (80)
for any $r\geq r_{0}$. For locally compact Polish spaces condition $0<\nu(C_{\psi}(r))<\infty$ is satisfied if $\psi$ has compact sub-levels sets $C_{\psi}(r)$ with non empty interior and $\nu$ is a Radon measure of full support. Also note that the l.h.s. minorization condition (80) is satisfied when $(x,y)\in(\mathbb{X}\times\mathbb{Y})^{\circ}\mapsto q(x,y)$ is a continuous positive function on the interior $(\mathbb{X}\times\mathbb{Y})^{\circ}:=(\mathbb{X}^{\circ}\times\mathbb{Y}^{\circ})$ of $(\mathbb{X}\times\mathbb{Y})$. Also note that
[…]
Indeed, if the inequality above holds then for any $0<\epsilon<\|\Theta\|$, the set $A_{\epsilon}:=\{\Theta\geq\epsilon\}=C_{1/\Theta}(1/\epsilon)$ is a non empty compact subset and we have
${\cal K}(\psi)\leq\epsilon~{}1_{S-A_{\epsilon}}\varphi+1_{A_{\epsilon}}(\Theta~{}\varphi)\leq\epsilon~{}\varphi+c_{\epsilon}\quad\mbox{\rm with}\quad c_{\epsilon}:=\|\Theta\|\sup_{A_{\epsilon}}\varphi.$ (81)
Theorem C.1.
Assume Condition ${\cal L}_{\varphi,\psi}({\cal K})$ is satisfied for some uniformly positive functions $(\varphi,\psi)\in({\cal B}(\mathbb{X})\times{\cal B}(\mathbb{Y}))$. Then, there exist constants $\kappa\geq 0$ and $\rho\in[0,1[$ that only depend on the parameters $(\epsilon,c)$ in (75) and on the function $\varepsilon(r)$ in (76) and such that
[…]
Theorem C.1 is a consequence of Lemma 2.3 in [36], see also Theorem 8.2.21 in [38] and Lemma 2.1 in [6]. Several examples satisfying the Lyapunov condition (75) and the local minorization conditions (82) are discussed in [6, 38].
Proof of Theorem 4.2
Assume that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$. From (39), for any $n>n_{0}$ we have the estimates
${\cal S}^{\circ}_{2n}(\psi)$ $=$ ${\cal S}_{2n-1}({\cal S}_{2n}(\psi))\leq\epsilon~{}{\cal S}_{2n-1}(\varphi)+c\leq\epsilon^{2}~{}\psi+(1+\epsilon)~{}c\quad\text{and}$
${\cal S}^{\circ}_{2n+1}(\varphi)$ $=$ ${\cal S}_{2n}({\cal S}_{2n+1}(\varphi))\leq\epsilon~{}{\cal S}_{2n}(\psi)+c\leq\epsilon^{2}~{}\varphi+(1+\epsilon)~{}c.$ (83)
Also note that for any $n>n_{0}$ the inequality (40) in Condition ${\cal D}_{\varphi,\psi}({\cal S})$ implies that
$\|((\delta_{y_{1}}-\delta_{y_{2}}){\cal S}_{2n-1}){\cal S}_{2n}\|_{\sf tv}\vee\|((\delta_{x_{1}}-\delta_{x_{2}}){\cal S}_{2n}){\cal S}_{2n+1}\|_{\sf tv}\leq 1-\varepsilon(r).$ (84)
The estimates in (83) imply that (33) holds, while the inequality (84) implies that (34) is satisfied.
Moreover, combining the fixed point equations in (13) with (83) we arrive at
$\sup_{n\geq n_{0}}\left(\|{\cal S}_{2n}(\psi)\|_{\varphi}\vee\|{\cal S}_{2n+1}(\varphi)\|_{\psi}\right)<\infty\quad\mbox{\rm and}\quad(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))\leq\frac{c}{1-\epsilon},$ (85)
and (39) together with (85) yields
$\sup_{n\geq n_{0}}(\pi_{2n}(\psi)\vee\pi_{2n+1}(\varphi))\leq c_{1}~{}(\lambda_{U}(\varphi)\vee\nu_{V}(\psi))$
for some $c_{1}>0$. Therefore, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied.
Finally, we observe that (39) ensures that inequality (75) is satisfied for the Markov transitions ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$, while (40) yields
$\chi_{\varphi}(r,{\cal S}_{2n})\vee\chi_{\psi}(r,{\cal S}_{2n+1})\leq 1-\varepsilon(r)$
and, hence, the inequality (76) also holds for ${\cal S}_{2n}$ and ${\cal S}_{2n+1}$. As a consequence, Theorem C.1 can be applied to show that there are parameters $\kappa\geq 0$ and $\rho\in[0,1[$ such that
[…]
Appendix D Schrödinger potential functions
D.1 Some quantitative estimates
Proposition D.1.
When $\jmath({\cal Q})>0$, for any $n\geq 0$ we have the uniform estimates
$\|\frac{d\pi_{2n}}{d\nu_{V}}\|\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|\leq\hbar({\cal Q})^{-2}\quad\mbox{and}\quad\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|\leq\hbar({\cal Q})^{-2}.$ (86)
Proof.
Recalling that $\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})$, for any given $x\in\mathbb{X}$ we have
$\hbar({\cal Q})~{}{\cal S}_{2n}(x,dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})=({\cal S}_{2n-1}{\cal S}_{2n})(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})}~{}{\cal S}_{2n}(x,dy_{2}).$
This yields, for any $(y_{1},\overline{y}_{1})\in\mathbb{Y}^{2}$, the estimate
$\hbar({\cal Q})^{2}~{}{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{}{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})$
and integrating w.r.t. $\nu_{V}(d\overline{y}_{1})$ and $\pi_{2(n-1)}(dy_{1})$ we find that
$\hbar({\cal Q})^{2}~{}\nu_{V}(dy_{2})\leq\pi_{2n}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{}\nu_{V}(dy_{2}).$
Similarly, integrating w.r.t. $\nu_{V}(dy_{1})$ and $\pi_{2(n-1)}(d\overline{y}_{1})$ we also obtain
$\hbar({\cal Q})^{2}~{}\pi_{2n}(dy_{2})\leq\nu_{V}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{}\pi_{2n}(dy_{2}).$
From the inequalities above we conclude that
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq\hbar({\cal Q})^{-2}\quad\text{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq\hbar({\cal Q})^{-2}.$
More generally these Radon Nikodym functions are unbounded and we need to carry a more refined analysis.
Proposition D.2.
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds for some $n_{0}\geq 1$ then there exists $c_{\epsilon}>0$ such that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$ (87)
Proof.
Using (15) and (33), we see that, for any $n\geq n_{0}$,
$\frac{d\pi_{2n}}{d\nu_{V}}$$\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~{}\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2n_{0}}\right)\left(\psi\right)$
$\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\left(\epsilon^{n-n_{0}}\psi+\frac{c}{1-\epsilon}\right)\leq c_{\epsilon}~{}\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~{}\psi,\quad\text{with}~{}c_{\epsilon}:=\left(1+\frac{c}{1-\epsilon}\right).$
This implies that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\text{and, similarly,}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$
Proposition D.3.
In addition, when Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, there exists $c_{\epsilon}>0$ such that
$\sup_{n\geq n_{0}}\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)\leq c_{\epsilon}~{}\left(\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}\right),$
as well as
$\sup_{n\geq n_{0}}\left(\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\right)\leq c_{\epsilon}~{}\left(\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\right).$
Proof.
If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$ then, by Theorem 4.2, the drift condition (33) also holds for any $n>n_{0}$, which implies that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}}}{d\nu_{V}}\|_{\psi}\quad\text{and, similarly,}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\pi_{2n_{0}+1}}{d\lambda_{U}}\|_{\varphi}.$
Arguing as in the proof of Proposition D.2, now combining the drift conditions (39) with (16), we find that
$\sup_{n\geq n_{0}}\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\leq c_{\epsilon}~{}\|\frac{d\nu_{V}}{d\pi_{2n_{0}}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\leq c_{\epsilon}~{}\|\frac{d\lambda_{U}}{d\pi_{2n_{0}+1}}\|_{\varphi}.$
Under the stronger Condition ${\cal H}_{\delta}(U,V)$, more refined estimates can be obtained. With that end, consider the functions
${\cal W}_{V,U}(x):=\log{{\cal Q}_{V}(e^{W_{U}})(x)}\quad\mbox{\rm and}\quad{\cal W}^{U,V}(y):=\log{{\cal R}_{U}(e^{W^{V}})(y)},$ (88)
with the integrated cost functions $(W_{U},W^{V})$ defined in (41) and the integral operators
${\cal Q}_{V}(x,dy)=e^{-W(x,y)}~{}\nu_{V}(dy)\quad\mbox{\rm and}\quad{\cal R}_{U}(y,dx)=e^{-W(x,y)}~{}\lambda_{U}(dx).$
Next technical lemma is pivotal.
Lemma D.4.
For any $n\geq 0$ we have
$\begin{array}[]{rcl}\log{{\cal Q}(\exp{(-V_{2n})})}&\geq&-W^{V}+\nu_{V}(V)\quad\text{and}\\ &&\\ -\nu_{V}(V)+{\cal W}^{U,V}\geq\log{{\cal R}(\exp{(-U_{2n})})}&\geq&-W_{U},\end{array}$ (89)
as well as
$U_{2n}\geq U-W^{V}+\nu_{V}(V).$
In addition, for any $n\geq 1$ we have
${\cal W}_{V,U}\geq\log{{\cal Q}(\exp{(-V_{2n})})}\geq-W^{V}+\nu_{V}(V)$ (90)
as well as
$\begin{array}[]{rcccl}U-W^{V}+\nu_{V}(V)&\leq&U_{2n}&\leq&U+{\cal W}_{V,U}\quad\text{and}\\ &&&&\\ V-W_{U}&\leq&V_{2n}&\leq&V+{\cal W}^{U,V}-\nu_{V}(V).\end{array}$ (91)
Proof.
Formula (10) yields the following monotone properties
$\begin{array}[]{rcccccl}\nu_{V}(V_{2(n+1)})&\leq&\nu_{V}(V_{2n})&=&\nu_{V}(V_{2(n+1)})+\mbox{\rm Ent}(\nu_{V}~{}|~{}\pi_{2n})&\leq&\nu_{V}(V_{0})=0,\\ &&&&&&\\ \lambda_{U}(U_{2(n+1)})&\leq&\lambda_{U}(U_{2n})&=&\lambda_{U}(U_{2(n+1)})+\mbox{\rm Ent}(\lambda_{U}~{}|~{}\pi_{2n+1})&\leq&\lambda_{U}(U_{0}).\end{array}$ (92)
Applying Jensen’s inequality, for any $n\geq 0$ we have
$\log{{\cal Q}(\exp{(-V_{2n})})(x)}$ $=$ $\log{\int~{}\nu_{V}(dy)~{}\exp{(-W(x,y)+V(y)-V_{2n}(y))}}$
$\geq$ $-W^{V}(x)+\nu_{V}(V)-\nu_{V}(V_{2n})\geq-W^{V}(x)+\nu_{V}(V).$
By (92) we also have $\lambda_{U}(U)\geq\lambda_{U}(U_{2n})$ and, therefore, for any $n\geq 0$ we obtain
$\log{{\cal R}(\exp{(-U_{2n})})(y)}$ $=$ $\log{\int~{}\lambda_{U}(dx)~{}\exp{(-W(x,y)+U(x)-U_{2n}(x))}}$
$\geq$ $-W_{U}(y)+\lambda_{U}(U)-\lambda_{U}(U_{2n})\geq-W_{U}(y)$
using Jensen’s inequality again. This ends the proof of (89).
The integral equations in (7) imply that for any $n\geq 0$ we have
$U_{2n}\geq U-W^{V}+\nu_{V}(V)\quad\mbox{\rm and}\quad V_{2(n+1)}\geq V-W_{U}.$
Using again (7) we see that
$U_{2(n+1)}-U=\log{{\cal Q}(e^{-V_{2(n+1)}})}$$\leq$ $\log{{\cal Q}_{V}(e^{W_{U}})}={\cal W}_{V,U}\quad\text{and}$
$V_{2(n+1)}-V=\log{{\cal R}(e^{-U_{2n}})}$$\leq$ $-\nu_{V}(V)+\log{{\cal R}_{U}(e^{W^{V}})}=-\nu_{V}(V)+{\cal W}^{U,V}.$
Proposition D.5.
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then
$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}\vee\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty.$ (93)
In addition, the exists some constant $c>0$ such that, for any $n\geq 1$,
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq c~{}e^{W_{U}}\quad\mbox{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq c~{}e^{W^{V}}.$ (94)
Proof.
Observe that
$\text{if}\quad\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty\quad\text{then}\quad e^{\delta U}=\frac{1}{e^{-\delta U}}\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}>0$
and this yields
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)=\lambda_{U}\left(e^{W^{V}}~{}e^{\delta U}\right)\geq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~{}\lambda_{U}\left(e^{W^{V}}\right).$
In the same vein, we obtain
$\nu_{(1-\delta)V}\left(e^{W_{U}}\right)=\nu_{V}\left(e^{W_{U}}~{}e^{\delta V}\right)\geq\|e^{-\delta V}\|_{\mathbb{L}_{\infty}(\nu)}^{-1}~{}\nu_{V}\left(e^{W_{U}}\right)$
and conclude that
$\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}$ $=$ $\|{\cal Q}_{V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{V}\left(e^{W_{U}}\right)$
$\leq$ $\|e^{-\delta V}\|_{\mathbb{L}_{\infty}(\nu)}~{}\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
as well as
$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{U}(e^{W^{V}})\|\leq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~{}\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$ by Condition ${\cal H}_{\delta}(U,V)$. This ends the proof of (93).
Using (91) we see that, for any $n\geq 1$,
$-{\cal W}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\cal W}_{V,U}+W^{V}-\nu_{V}(V)$
as well as
$-{\cal W}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\cal W}^{U,V}+W_{U}-\nu_{V}(V).$
Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (99) and (100) in the proof of the commutation formulae (11), we readily obtain the estimates
$e^{\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~{}e^{-W_{U}}\leq\frac{d\pi_{2n}}{d\nu_{V}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}~{}e^{W_{U}}$
as well as
$e^{\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~{}e^{-W^{V}}\leq\frac{d\pi_{2n+1}}{d\lambda_{U}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}~{}e^{W^{V}}$
Remark D.6.
The functions $({\cal W}_{V,U},{\cal W}^{U,V})$ defined in (88) can be computed explicitly for the the Gaussian model defined in (50) and (51).
We set $\sigma_{\beta}=\beta\sigma\beta^{\prime}$ and $\beta_{\tau}:=\tau^{-1}\beta$. With this notation at hand, after some elementary computations (cf. for instance Exemple 6.6 in [2]) we have
${\cal W}_{V,U}(x)=\frac{1}{2}~{}\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~{}(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~{}(\beta_{\tau}~{}(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
and
${\cal W}^{U,V}(y)=\frac{1}{2}~{}\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\frac{1}{2}(y-\overline{m})^{\prime}\left(\beta_{\tau}\,\sigma\,\beta_{\tau}^{\prime}-\tau^{-1}\right)(y-\overline{m})-(y-\overline{m})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$ (95)
Note that one readily obtain the relationships
[…]
Using (94), Condition ${\cal H}_{\delta}(U,V)$ ensures that for any $n\geq 1$ we have
$e^{-\delta V}~{}\left(\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\right)\leq c~{}e^{-{\cal V}_{\delta}}\quad\mbox{\rm and}\quad e^{-\delta U}~{}~{}\left(\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\right)\leq c~{}e^{-{\cal U}_{\delta}}$
for some constant $c<\infty$. These estimates immediately yield the following result.
Proposition D.7.
Under the assumptions of Proposition D.5, for any $n\geq 1$ we have
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\in{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\in{\cal B}_{\varphi}(\mathbb{X}).$
In addition, we have the uniform estimates
$\sup_{n\geq 1}~{}\left(\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|_{\psi}\right)<\infty\quad\mbox{and}\quad\sup_{n\geq 1}~{}\left(\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|_{\varphi}\right)<\infty.$
D.2 Lyapunov inequalities
The result below is a fairly direct consequence of the estimates presented in Lemma D.4.
Lemma D.8.
For any $n\geq 1$ we have the estimates
$e^{\nu_{V}(V)}~{}e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{x}{\cal S}_{2n}}{d\delta_{x}{\cal Q}_{V}}(y)\leq e^{-\nu_{V}(V)}~{}e^{W^{V}(x)+W_{U}(y)}$
and
$e^{\nu_{V}(V)}~{}e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}\leq\frac{d\delta_{y}{\cal S}_{2n+1}}{d\delta_{y}{\cal R}_{U}}(x)\leq e^{-\nu_{V}(V)}~{}e^{W^{V}(x)+W_{U}(y)}.$
Proof.
Using the Sinkhorn transition operators as expressed in (9) together with (90) and (91), we find that, for any $n\geq 1$,
${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{(V-V_{2n})(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\geq e^{\nu_{V}(V)}e^{-({\cal W}_{V,U}(x)+{\cal W}^{V,U}(y))}{\cal Q}_{V}(x,dy)~{}$
and, in the same vein,
${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~{}e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\geq e^{\nu_{V}(V)}~{}e^{-({\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{}{\cal R}_{U}(y,dx).$
Similarly, one can resort to (90) and (91) again to obtain the lower bounds
${\cal S}_{2n}(x,dy)$ $=$ $\frac{{\cal Q}_{V}(x,dy)e^{V-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\leq e^{-\nu_{V}(V)}~{}e^{W^{V}(x)}~{}{\cal Q}_{V}(x,dy)~{}e^{W_{U}(y)}$
and, in the same way,
${\cal S}_{2n+1}(y,dx)$ $=$ $\frac{{\cal R}_{U}(y,dx)~{}e^{(U-U_{2n})(x)}}{{\cal R}(e^{-U_{2n}})(y)}\leq e^{-\nu_{V}(V)}~{}e^{W_{U}(y)}~{}{\cal R}_{U}(y,dx)~{}e^{W^{V}(x)}.$
With these estimates at hand, we can now state and prove the following Lyapunov inequalities.
Proposition D.9.
Assume that Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. Then, the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$ satisfy the Lyapunov inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi\leq c_{\delta}~{}e^{-{\cal U}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{X})\quad\mbox{and}\quad{\cal S}_{2n+1}\left(\varphi\right)/\psi\leq c_{\delta}~{}e^{-{\cal V}_{\delta}}~{}\in{\cal B}_{0}(\mathbb{Y})$ (96)
for every $n\geq 1$ and some finite constant $c_{\delta}<\infty$.
Proof.
By Lemma D.8 we have
${\cal S}_{2n}(x,dy)~{}\psi(y)/\varphi(x)$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal U}_{\delta}(x)}~{}{\cal Q}_{(1-\delta)V}(x,dy)~{}e^{W_{U}(y)},\quad\text{and}$
${\cal S}_{2n+1}(y,dx)~{}\varphi(x)/\psi(y)$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal V}_{\delta}(y)}~{}{\cal R}_{(1-\delta)U}(y,dx)~{}e^{W^{V}(x)},$
which (via (88)) imply the inequalities
${\cal S}_{2n}\left(\psi\right)/\varphi$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal U}_{\delta}}~{}e^{{\cal W}_{(1-\delta)V,U}},\quad\text{and}$
${\cal S}_{2n+1}\left(\varphi\right)/\psi$$\leq$ $e^{-\nu_{V}(V)}~{}e^{-{\cal V}_{\delta}}~{}e^{{\cal W}^{(1-\delta)U,V}}.$
On the other hand, we have the estimates
$\|e^{{\cal W}_{(1-\delta)V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}=\|{\cal Q}_{(1-\delta)V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
and
$\|e^{{\cal W}^{(1-\delta)U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{(1-\delta)U}(e^{W^{V}})\|_{\mathbb{L}_{\infty}(\nu)}\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ when Condition ${\cal H}_{\delta}(U,V)$ holds. This yields (96) with the constant
$c_{\delta}:=e^{-\nu_{V}(V)}~{}\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\left(\nu_{(1-\delta)V}\left(e^{W_{U}}\right)\vee\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\right).$
D.3 Proof of Theorem 4.3
Following exactly the same argument leading to (81) in Appendix C, the Lyapunov inequalities (96) in Proposition D.9 ensure that the drift conditions stated in (39) hold with with $n_{0}=1$ and $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$.
For $\lambda$-almost any $x\in C_{\varphi}(r)$ we have
$e^{{\cal W}_{V,U}(x)}={\cal Q}_{V}(e^{W_{U}})(x)$ $=$ $\int~{}e^{W_{U}(y)-W(x,y)}~{}\nu_{V}(dy)$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\int~{}e^{W_{U}(y)}~{}\nu_{V}(dy)\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)$
and we conclude that
$w_{V,U}(r):=\sup_{C_{\varphi}(r)}{\cal W}_{V,U}<\infty.$
Similarly, for $\nu$-almost any $y\in C_{\psi}(r)$ we have
$e^{{\cal W}^{U,V}(y)}={\cal R}_{U}(e^{W^{V}})(y)$ $=$ $\int~{}e^{-W(x,y)+W^{V}(x)}~{}\lambda_{U}(dx)$
$\leq$ $\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{U}(e^{W^{V}})\leq\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~{}\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$
which yields
$w^{U,V}(r):=\sup_{C_{\psi}(r)}{\cal W}^{U,V}<\infty.$
On the other hand, using Lemma D.8 for any $x\in C_{\varphi}(r)$ we arrive at
${\cal S}_{2n}(x,dy)$$\geq$ $e^{\nu_{V}(V)}~{}e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{}\nu_{V}(dy)~{}1_{C_{\psi}(r)}(y)$
$\geq$ $e^{-w(r)}~{}\nu_{V}\left(C_{\psi}(r)\right)~{}{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad{\cal Y}_{r}(dy):=\frac{\nu_{V}(dy)~{}1_{C_{\psi}(r)}(y)}{\nu_{V}\left(C_{\psi}(r)\right)}$
and the parameter
$w(r):=\sup_{C_{\varphi}(r)\times C_{\psi}(r)}W+w_{V,U}(r)+w^{U,V}(r)-\nu_{V}(V).$
Note that for any $r>1$ we have
$\nu_{V}\left(C_{\psi}(r)\right)$ $=$ $e^{-V_{\star}}~{}\nu\left(e^{-(V-V_{\star})}~{}1_{(V-V_{\star})\leq\delta^{-1}\log{r}}\right)$
$\geq$ $w_{0}(r):=e^{-V_{\star}}~{}e^{-(\delta^{-1}\log{r})}~{}\nu\left(C_{\psi}(r)\right)>0$
and this yields, for any $x\in C_{\varphi}(r)$, the estimate
${\cal S}_{2n}(x,dy)\geq\varepsilon_{0}(r)~{}{\cal Y}_{r}(dy)\quad\mbox{\rm with}\quad\varepsilon_{0}(r):=w_{0}(r)~{}e^{-w(r)}.$
In the same vein, for any $y\in C_{\psi}(r)$ we have
${\cal S}_{2n+1}(y,dx)$$\geq$ $e^{\nu_{V}(V)}~{}e^{-(W(x,y)+{\cal W}_{V,U}(x)+{\cal W}^{U,V}(y))}~{}\lambda_{U}(dx)~{}1_{C_{\varphi}(r)}(x)$
and, arguing as above, we arrive at
$\lambda_{U}\left(C_{\varphi}(r)\right)\geq w_{1}(r):=e^{-U_{\star}}~{}e^{-(\delta^{-1}\log{r})}~{}\lambda\left(C_{\varphi}(r)\right)>0.$
Thus, for any $y\in C_{\psi}(r)$ we have
${\cal S}_{2n+1}(y,dx)\geq\varepsilon_{1}(r)~{}{\cal X}_{r}(dx)$
where
$\varepsilon_{1}(r):=w_{1}(r)~{}e^{-w(r)}\quad\mbox{\rm and}\quad{\cal X}_{r}(dx):=\frac{\lambda_{U}(dx)~{}1_{C_{\varphi}(r)}(x)}{\lambda_{U}\left(C_{\varphi}(r)\right)}.$
The above estimates ensure that the local minorization condition (40) is met with $n_{0}=1$ and the function
$\varepsilon(r):=\varepsilon_{0}(r)\wedge\varepsilon_{1}(r)>0.$
Hence, we conclude that Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied with $n_{0}=1$.
Using (13) and (15), for any $n\geq p\geq 1$ we verify that
$\frac{d\pi_{2n}}{d\nu_{V}}(y)-1=\left[(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right]\left(\frac{d\pi_{2p}}{d\nu_{V}}\right),$
which implies the inequality
$\left|\frac{d\pi_{2n}}{d\nu_{V}}(y)-1\right|\leq\|\frac{d\pi_{2p}}{d\nu_{V}}\|_{\psi}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}.$
On the other hand, combining Proposition D.7 with the norm equivalence (19) and Theorem 4.2, we see that there exists some $\kappa>0$ and some constants $c_{\kappa,i}>0$, $i=1,2$, such that
$~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi}$$\leq$ $c_{\kappa,1}~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\delta_{y}-\nu_{V})\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2(p+1)}\right)\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\psi_{\kappa}}$
$\leq$ $c_{\kappa,1}~{}\rho^{2(n-p)}~{}|\delta_{y}-\nu_{V}|_{\psi_{\kappa}}\leq c_{\kappa,2}~{}\rho^{2(n-p)}~{}(\psi(y)+\nu_{V}(\psi)).$
Therefore, for any $n\geq 1$, we obtain
$\left\|\frac{d\pi_{2n}}{d\nu_{V}}-1\right\|_{\psi}\leq c_{\kappa,3}~{}\rho^{2n}~{}\quad\text{and, similarly,}\quad\left\|\frac{d\nu_{V}}{d\pi_{2n}}-1\right\|_{\psi}\leq c_{\kappa,4}~{}\rho^{2n},$ (97)
for some constants $c_{\kappa,i}$, $i=3,4$. By the same argument, using (13) we have, for any $n\geq p\geq 1$,
$\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)-1=\left[(\delta_{x}-\lambda_{U})({\cal S}^{\circ}_{2n+1}\ldots{\cal S}^{\circ}_{2p+1})\right]\left(\frac{d\pi_{2p-1}}{d\lambda_{U}}\right)$
and arrive at
$\left\|\frac{d\pi_{2n+1}}{d\lambda_{U}}-1\right\|_{\varphi}\vee\left\|\frac{d\lambda_{U}}{d\pi_{2n+1}}-1\right\|_{\varphi}\leq c_{\kappa,5}~{}\rho^{2n}~{}$ (98)
for any $n\geq 1$ and some constant $c_{\kappa,5}$.
Finally, note that for any function $f(x)>0$ we have
$\left|\log{f(x)}\right|$ $=$ $\log{\left(1+(f(x)-1)\right)}~{}1_{f(x)\geq 1}+\log{\left(1+\frac{1}{f(x)}-1\right)}~{}1_{{1}/{f(x)}\geq 1}$
$\leq$ $\left(f(x)-1\right)~{}1_{f(x)\geq 1}+\left(\frac{1}{f(x)}-1\right)~{}1_{{1}/{f(x)}\geq 1},$
where the last assertion comes from the fact that $\log{(1+u)}\leq u$, for any $01+u>$ 0. This implies that
$\|\log{f}\|_{\varphi}=\|\log{1/f}\|_{\varphi}\leq\|f-1\|_{\varphi}+\|1/f-1\|_{\varphi}$
The inequality above combined with (97) and (98) yields the estimates
$\left\|\log{\frac{d\pi_{2n}}{d\nu_{V}}}\right\|_{\psi}\vee\left\|\log{\frac{d\pi_{2n+1}}{d\lambda_{U}}}\right\|_{\varphi}\leq c_{6,\kappa}~{}\rho^{2n},$
for some finite constant $c_{\kappa,6}$. This completes the proof of (42).
Appendix E Some technical proofs
Proof of (11) and (12)
We first note, using (5) and (6), that
$\frac{d{\cal P}_{2n}}{d{\cal P}_{2n+1}}(x,y)=\frac{d\pi_{2n}}{d\nu_{V}}(y)=\exp{\left(V_{2n+1}(y)-V_{2n}(y)\right)}$ (99)
and, in the same vein,
$\frac{d{\cal P}_{2n+1}}{d{\cal P}_{2(n+1)}}(x,y)=\frac{d\pi_{2n+1}}{d\lambda_{U}}(x)=\exp{\left(U_{2(n+1)}(x)-U_{2n+1}(x)\right)}.$ (100)
If we additionally observe that
${\cal P}_{2(n+1)}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{}{\cal P}_{2n}(d(x,y))~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\cal P}_{2n+1}(d(x,y))$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n-1}}(x)~{}{\cal P}_{2n-1}(d(x,y))~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)$
then, using (4), we obtain the formulae
${\cal S}_{2(n+1)}(x,dy)$ $=$ $\frac{d\lambda_{U}}{d\pi_{2n+1}}(x)~{}{\cal S}_{2n}(x,dy)~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)\quad\text{and}$
${\cal S}_{2n+1}(y,dx)$ $=$ $~{}\frac{d\nu_{V}}{d\pi_{2n}}(y)~{}{\cal S}_{2n-1}(y,dx)~{}\frac{d\lambda_{U}}{d\pi_{2n-1}}(x),$
and the commutation properties now follow from elementary manipulations.
Proof of (37)
Assume Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied. Combining (33) with (13) we arrive at
$\nu_{V}(\psi)=\nu_{V}({\cal S}_{2n}^{\circ}(\psi))\leq\epsilon~{}\nu_{V}(\psi)+c\quad\text{which implies}\quad\nu_{V}(\psi)\leq\frac{c}{1-\epsilon}.$
In a similar way, one can also show that
$\lambda_{U}(\varphi)\leq\frac{c}{1-\epsilon}.$
Using the Markov evolution equations in (14) together with (33) we find that, for any $n\geq n_{0}$,
$\pi_{2n}(\psi)$ $=$ $\pi_{2(n-1)}({\cal S}_{2n}^{\circ}(\psi))$
$\leq$ $\epsilon~{}\pi_{2(n-1)}(\psi)+c\leq\epsilon^{n-n_{0}}~{}\pi_{2n_{0}}(\psi)+\frac{c}{1-\epsilon}$
and, by the same argument,
$\pi_{2n+1}(\varphi)$ $=$ $\pi_{2n-1}({\cal S}_{2n+1}^{\circ}(\varphi))$
$\leq$ $\epsilon~{}\pi_{2n-1}(\varphi)+c\leq\epsilon^{n-n_{0}}~{}\pi_{2n_{0}+1}(\varphi)+\frac{c}{1-\epsilon}.$
Proof of Proposition 5.4
It is straightforward to show that
$2W_{U}(y)$ $=$ $\left\|\tau^{-1/2}(y-m_{0})\right\|_{F}^{2}+\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\log{\mbox{\rm det}(2\pi\tau)}\quad\text{and}$
$2W^{V}(x)$ $=$ $\left\|\tau^{-1/2}(\beta(x-m)+(m_{0}-\overline{m}))\right\|_{F}^{2}+~{}\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\log{\mbox{\rm det}(2\pi\tau)},$
with $m_{0}:=(\alpha+\beta m)$ and $\sigma_{\beta}:=(\beta\sigma\beta^{\prime})$, hence $W_{U}$ and $W^{V}$ are locally bounded. Moreover, for any $\delta\geq 0$ we obtain
$2(\delta U-W^{V})$ $=$ $-\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\delta~{}\log{\mbox{\rm det}(2\pi\sigma)}-\log{\mbox{\rm det}(2\pi\tau)}$
$+(x-m)^{\prime}\left(\delta~{}\sigma^{-1}-\beta^{\prime}\tau^{-1}\beta\right)(x-m)$
$-2(m_{0}-\overline{m})^{\prime}\tau^{-1}\beta(x-m)-(m_{0}-\overline{m})^{\prime}\tau^{-1}(m_{0}-\overline{m})$
and, in the same way,
$2(\delta V-W_{U})$ $=$ $-\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\delta~{}\log{\mbox{\rm det}(2\pi\overline{\sigma})}-\log{\mbox{\rm det}(2\pi\tau)}$
$+(y-\overline{m})^{\prime}\left(\delta~{}\overline{\sigma}^{-1}-\tau^{-1}\right)(y-\overline{m})$
$-2(\overline{m}-m_{0})^{\prime}\tau^{-1}(y-\overline{m})-(\overline{m}-m_{0})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$
Therefore,
${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{R}^{d})\quad\mbox{\rm and}\quad{\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{R}^{d})$
if, and only if,
$\beta^{\prime}\tau^{-1}\beta<\delta~{}\sigma^{-1}\quad\mbox{\rm and}\quad\tau^{-1}<\delta~{}\overline{\sigma}^{-1}.$
On the other hand, straightforward manipulations yield
$2(W^{V}-(1-\delta)U)$ $=$ $\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})-(1-\delta)~{}\log{\mbox{\rm det}(2\pi\sigma)}+\log{\mbox{\rm det}(2\pi\tau)}$
$-(x-m)^{\prime}\left((1-\delta)~{}\sigma^{-1}-\beta^{\prime}\tau^{-1}\beta\right)(x-m)$
$+2(m_{0}-\overline{m})^{\prime}\tau^{-1}\beta(x-m)+(m_{0}-\overline{m})^{\prime}\tau^{-1}(m_{0}-\overline{m}),$
which shows that
$\lambda_{(1-\delta)U}(e^{W^{V}})<\infty\Longleftrightarrow\beta^{\prime}\tau^{-1}\beta<(1-\delta)~{}\sigma^{-1}.$
In the same vein, we have
$2(W_{U}-(1-\delta)V)$ $=$ $\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})-(1-\delta)~{}\log{\mbox{\rm det}(2\pi\overline{\sigma})}-\log{\mbox{\rm det}(2\pi\tau)}$
$-(y-\overline{m})^{\prime}\left((1-\delta)~{}\overline{\sigma}^{-1}-\tau^{-1}\right)(y-\overline{m})$
$+2(\overline{m}-m_{0})^{\prime}\tau^{-1}(y-\overline{m})+(\overline{m}-m_{0})^{\prime}\tau^{-1}(\overline{m}-m_{0}),$
which shows that
$\nu_{(1-\delta)V}(e^{W_{U}})<\infty\Longleftrightarrow\tau^{-1}<(1-\delta)~{}\overline{\sigma}^{-1}.$
Proof of Proposition 5.5
Let us first recall that
$\lambda_{U}(dx)=\sum_{i\in E_{0}}\xi_{0}(i)~{}\lambda_{U_{i}}(dx)\quad\mbox{\rm and}\quad\nu_{V}(dy)=\sum_{j\in E_{1}}\xi_{1}(j)~{}\nu_{V_{j}}(dx),$
which implies that
$W_{U}=\sum_{i\in E_{0}}~{}\xi_{0}(i)~{}W_{U_{i}}\quad\mbox{\rm and}\quad W^{V}=\sum_{j\in E_{1}}~{}\xi_{1}(j)~{}W^{V_{j}}$ (101)
and ensures that the functions $W_{U}$ and $W^{V}$ are locally bounded.
Moreover, we note that
$\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}\leq e^{-\inf_{i\in E_{0}}U_{i}},$
which implies that
$-U=\log{\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}}}\leq-\inf_{i\in E_{0}}U_{i}$
and, hence,
$U\geq\inf_{i\in E_{0}}U_{i}.$
As a consequence, we readily obtain the inequalities
$\delta U-W^{V}\geq\inf_{i\in E_{0}}\sum_{j\in E_{1}}~{}\xi_{1}(j)~{}(\delta~{}U_{i}-~{}W^{V_{j}})\geq\inf_{(i,j)\in(E_{0}\times E_{1})}~{}(\delta~{}U_{i}-~{}W^{V_{j}})$
which, in turn, enable us to show that
$\left\{x~{}:~{}\delta U(x)-W^{V}(x)\leq r\right\}$$\subset$ $\left\{x~{}:~{}\inf_{(i,j)\in(E_{0}\times E_{1})}~{}(\delta~{}U_{i}(x)-~{}W^{V_{j}}(x))\leq r\right\}$
$=$ $\bigcup_{(i,j)\in(E_{0}\times E_{1})}\left\{x~{}:~{}(\delta~{}U_{i}(x)-~{}W^{V_{j}}(x))\leq r\right\}$
and this ensures that ${\cal U}_{\delta}=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$. In the same vein, one can also prove that ${\cal V}_{\delta}=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$.
On the other hand, using the fact that $(a+b)^{1-\delta}\leq a^{1-\delta}+b^{1-\delta}$ for any $a,b\geq 0$, we have
$\lambda_{(1-\delta)U}(dx)$ $=$ $\left(\sum_{i\in E_{0}}\xi_{0}(i)~{}e^{-U_{i}(x)}\right)^{1-\delta}~{}\lambda(dx)$ (102)
$\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}e^{-(1-\delta)U_{i}(x)}~{}\lambda(dx)=\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}\lambda_{(1-\delta)U_{i}}(dx)$
and, by way of Hölder’s inequality, we also have
$\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~{}e^{\xi_{1}(j)~{}W^{V_{j}}}\right)\leq\prod_{j\in E_{1}}~{}\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}.$ (103)
Combining (101), (102) and (103) we obtain the estimates
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$$\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}\lambda_{(1-\delta)U_{i}}\left(\prod_{j\in E_{1}}~{}e^{\xi_{1}(j)~{}W^{V_{j}}}\right)$
$\leq$ $\sum_{i\in E_{0}}\xi_{0}(i)^{1-\delta}~{}\prod_{j\in E_{1}}~{}\left(\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)\right)^{\xi_{1}(j)}$
from which we conclude that $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)<\infty$ (because $\lambda_{(1-\delta)U_{i}}\left(e^{W^{V_{j}}}\right)<\infty$ for almost every $(i,j)$). Using the same argument, we can also prove that $\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$.